### v2

# Constants of motion in gravitational self-force theory

Abstract
Synergies between self-force theory and other approaches to the gravitational two-body problem have traditionally relied on calculations of gauge-invariant observables as functions of orbital frequencies. However, in self-force theory one can also define a complete set of constants of motion: energy, azimuthal angular momentum, and radial and polar actions. Here we outline how directly utilizing these constants allows for more straightforward comparisons and hybridizations across the parameter space, as well as more streamlined waveform generation through flux-balance laws. Restricting to the case of eccentric, nonspinning binaries and first order in self-force, we compute the constants of motion and the corrections to fundamental frequencies numerically as well as analytically (to 9PN in a post-Newtonian expansion), establishing consistency with the highest-order (4PN) results available from post-Newtonian theory. We also apply the results to identify the perturbed locations of special curves in the parameter space: circular orbits and the separatrix between bound and plunging orbits.
I Introduction
Much of the progress in gravitational-wave source modeling has relied on combining multiple methods [102, 4], each of which is accurate in a different regime of the two-body problem: post-Newtonian (PN) theory, accurate in the early inspiral [38]; self-force (SF) theory, accurate for small mass ratios [17]; numerical relativity, best suited to comparable masses in the late inspiral, merger, and ringdown [21]; black hole perturbation theory, applicable to the final ringdown [27]; and most recently, post-Minkowskian (PM) theory, accurate for weak-field, unbound orbits [61]. Tools and results from these methods have been combined in a myriad of ways (e.g., [20, 10, 128, 140, 104, 8, 44, 58, 59, 150, 120, 161, 141, 5, 144, 82, 109, 135, 60, 16, 33, 32, 97, 31, 110, 89, 125, 169, 92, 88, 87]), often through the framework of effective one-body (EOB) theory [42, 43].
This synergism is enabled by calculations of the same quantities in multiple formalisms in order to translate information between them. Frequently the computed quantities are waveforms themselves, but in many cases the translation occurs at the level of orbital dynamics—or at the level of waveforms as functions of orbital parameters [165]. In PN and PM theory, the orbital dynamics is very often written in the language of Hamiltonian mechanics, with well-defined notions of conserved energy, angular momentum, and action variables [57, 26, 25, 73]. In self-force theory, such conserved quantities have been essential for the understanding and exploitation of the leading-order (0SF), geodesic motion of the secondary object (of mass $m_{2}$) around the primary black hole (of mass $m_{1}\geq m_{2}$[146, 86, 130]. They have also played a key role in modeling the effects of the secondary’s spin [171, 172, 147, 139, 136, 127, 81, 170, 113]. But beyond the test-body limit, when the secondary object perturbs the primary’s spacetime and experiences a self-force, such conserved quantities have been little explored.
Notions of conserved energy, angular momentum, and radial action at first subleading order in the mass ratio $\varepsilon:=m_{2}/m_{1}$ (1SF) were first proposed in Refs. [100, 103]. There they were effectively defined as the quantities satisfying the first law of binary mechanics [100, 79, 35, 101, 103, 76, 37, 138, 137, 77] for circular or eccentric orbits of the secondary around a Schwarzschild primary. Reference [76] expanded upon this approach and extended it to the case of generic bound orbits around a spinning, Kerr black hole, again defining the conserved quantities as the ones satisfying a first law of binary mechanics. However, this approach required an ad hoc “renormalization” of the conserved quantities. More recently, a complete Hamiltonian analysis [107] building on Refs. [76, 39] has shown that the “renormalized” variables precisely correspond to natural notions in Hamiltonian mechanics: a conserved energy that is the on-shell value of the Hamiltonian; and action variables—radial, polar, and azimuthal, the last of which represents azimuthal angular momentum—with the traditional meaning of invariant integrals over tori in the orbital phase space [11].
In this paper, we begin to explore the behavior and utility of these 1SF conserved quantities in more detail, specializing to the case of eccentric bound orbits around a nonspinning, Schwarzschild black hole. Access to these quantities brings several advantages:
  1. 1.
    It allows one to cast waveform generation in a streamlined form in terms of balance laws, in which the waveform’s evolution is dictated by the slow evolution of the conserved quantities, and that slow evolution is itself determined from the gravitational-wave fluxes to infinity and down the primary’s horizon. We expect this to be crucial at the first subleading order in $\varepsilon$—so-called first post-adiabatic (1PA) order [130, 113]. Such 1PA models are required for modeling extreme-mass-ratio inspirals (EMRIs) [4, 45], key sources for space-based detectors such as LISA. Currently, the only extant 1PA waveform models [167, 114], which have been restricted to quasicircular orbits, have relied on this balance-law approach. We expect their generalization to eccentric orbits to do so as well.
  2. 2.
    It facilitates communication between SF, PN, and PM theory. For example, formulations of waveform generation in terms of balance laws enable simple hybridizations at the level of conserved quantities and fluxes; see Refs. [88, 87]. As outlined in those references, hurdles in second-order self-force (2SF) calculations [124] mean that such a hybrid approach might be essential for the construction of 1PA waveform models across the binary parameter space in the near term. Direct applications to EOB models can also be envisioned.
  3. 3.
    It provides a simple way of computing the corrections to the binary’s fundamental frequencies, which have a central role in EMRI modeling [130, 91, 96, 167, 49, 113]. Traditionally, translation of 1SF results into other formalisms has relied on computations of invariant observables as functions of these frequencies [102, 6, 17].
  4. 4.
    Following the path laid out by Le Tiec [103], it provides a clean characterization of the orbital phase space, helping to identify special surfaces such as circular orbits and the separatrix between bound and plunging orbits.
To capitalize on these benefits, here we calculate the 1SF energy, angular momentum, radial action, and fundamental frequencies across the parameter space of bound orbits in Schwarzschild spacetime, using both numerical and analytical (post-Newtonian) methods [^1] In the context of gravitational self-force, post-Newtonian expansion can be obtained to very high orders, but are restricted to the leading or subleading order in the mass ratio; we refer to them as “analytical self force” [145] or “SF-PN”. Conversely, “traditional PN” methods [38] are those which can access all orders in the mass ratio; these are typically more involved and current literature is restricted to lower orders. from Refs. [121, 118]. We verify that the results agree with those of traditional PN theory at all available PN orders (through 4PN [152]), and we present several applications of them. We expect these results to be widely useful across the community, in the construction of 1PA SF models, in validation of other models, and in the development of EOB and other hybrid models. For ease of use, we also delineate the region where high-order analytical results (carried to 9PN in the weak-field expansion) become more accurate than our numerical data.
The paper is organized as follows. In Sec. II we review the SF formalism and the 1SF conserved quantities. In Sec. III we review the relevant equations at geodesic, 0SF order, including PN expansions of geodesic relationships. Section IV describes our calculation of the 1SF corrections, as well as the comparison between our numerical and analytical results. Section V applies our results to identify the location of special curves in the parameter space, as mentioned above, and outlines waveform generation schemes using our results. We conclude in Sec. VI with a discussion of future applications. Some technical details and lengthy expressions are relegated to appendices. We provide our data and PN expansions in the Supplemental Material [1].
We use geometric units with $G=c=1$ and the mostly positive signature (+++)(-+++). Repeated indices are summed over even if they have the same vertical placement, as in $k_{A}\varphi_{A}:=k_{r}\varphi_{r}+k_{\phi}\varphi_{\phi}$.
II Formalism
Our review of conserved quantities in SF theory broadly follows Ref. [107], though with some minor changes in notation that we note along the way. Rather than specializing to conservative dynamics, we emphasize the role of the conserved quantities in the full, dynamical, radiating spacetime, situating them within the broader conceptual framework of multiscale expansions [86, 117, 130, 116, 113, 168, 107] that underlies most modern SF calculations and enables rapid SF waveform generation [52, 96, 49].
II.1 Multiscale formulation of self-force theory
In gravitational SF theory, the spacetime metric is expanded in powers of the mass ratio $\varepsilon$,
$g_{\alpha\beta}=g^{(0)}_{\alpha\beta}+\varepsilon h^{(1)}_{\alpha\beta}+\varepsilon^{2}h^{(2)}_{\alpha\beta}+{\cal O}(\varepsilon^{3}),$ (1)
and the secondary is reduced to a point particle orbiting in the spacetime of the primary. Here we restrict to the special case of a Schwarzschild background metric $g^{(0)}_{\alpha\beta}$ of mass $m_{1}$. Assuming the secondary is also nonspinning, its motion is geodesic not in $g_{\alpha\beta}$ (which is singular at the particle’s position) or $g^{(0)}_{\alpha\beta}$ but in a certain effective metric [64, 66, 131, 133, 134, 84, 83]
$\tilde{g}_{\alpha\beta}=g^{(0)}_{\alpha\beta}+\varepsilon h^{{\rm R}(1)}_{\alpha\beta}+\varepsilon^{2}h^{{\rm R}(2)}_{\alpha\beta}+{\cal O}(\varepsilon^{3}),$ (2)
which is regular and satisfies the vacuum Einstein equation even at the particle’s position. The geodesic equation in $\tilde{g}_{\alpha\beta}$ can equivalently be written in terms of the background metric and perturbations as [132]
$\frac{D^{2}x^{\alpha}_{p}}{d\tau^{2}}=\varepsilon f^{\alpha}_{(1)}+\varepsilon^{2}f^{\alpha}_{(2)}+{\cal O}(\varepsilon^{3}).$ (3)
Here $x^{\alpha}_{p}(\tau)=(t_{p}(\tau),r_{p}(\tau),\theta_{p}(\tau),\phi_{p}(\tau))$ is the particle’s trajectory in Schwarzschild coordinates; $\tau$ is proper time measured in $g^{(0)}_{\alpha\beta}$; and $D/d\tau:=u^{\alpha}\nabla_{\alpha}$ is the Schwarzschild covariant derivative along the particle’s worldline, with $u^{\alpha}:=dx^{\alpha}_{p}/d\tau$. The first- and second-order self-forces per unit mass, $f^{\alpha}_{(1)}$ and $f^{\alpha}_{(2)}$, are constructed from $h^{{\rm R}(1)}_{\alpha\beta}$ and $h^{{\rm R}(2)}_{\alpha\beta}$. See Eqs. (39) and (40) of Ref. [132], for example; we do not display these formulas here because we advocate for the use of balance laws to avoid explicit calculations of the self-forces.
The multiscale expansion of the orbital dynamics (and of the spacetime metric) is based on reformulating the self-forced motion in terms of fast and slow variables that cleanly separate oscillatory effects from slow evolution. These variables are best understood as coordinates on the orbital phase space. To exploit the stationarity of the background, and ultimately to obtain waveforms in terms of the time of an asymptotic observer, we use Schwarzschild time $t$ as the parameter along the particle’s phase-space trajectory rather than proper time $\tau$. Since self-forced orbits in a Schwarzschild background are confined to a plane, we can also freely specify that they lie in the equatorial plane, $\theta=\pi/2$. With $t$ reduced to a parameter along the trajectory and with polar motion eliminated, the orbital phase space is reduced from eight dimensions to four, with phase-space coordinates given by the positions and momenta $(x^{A}_{p},p_{A})$, where $A=r,\phi$ and $p_{A}=m_{2}g_{AB}^{(0)}u^{B}$. [^2] Reference [107] instead uses momenta $\tilde{p}_{A}=m_{2}\tilde{g}_{A\beta}\frac{dx_{p}^{\beta}}{d\tilde{\tau}}$, which are canonically conjugate to $x^{A}_{p}$ at all SF orders in a certain pseudo-Hamiltonian sense. $p_{A}$ in that reference denotes yet another momentum, which is conjugate to $x^{A}_{p}$ in the 1SF Hamiltonian we review below. The $p_{A}$ we use here is conjugate to $x^{A}_{p}$ only with respect to the background-geodesic Hamiltonian.
Appropriate fast variables $\varphi_{A}$ and slow variables $\pi_{A}$ are found via an $\varepsilon$-dependent transformation $(x^{A}_{p},p_{A})\mapsto(\varphi_{A},\pi_{A})$, [^3] Our variables $(\varphi_{A},\pi_{A})$ are denoted $(\mathring{\varphi}^{A},\mathring{\pi}_{A})$ in Ref. [107], $(\varphi_{A},p_{\varphi}^{A})$ in Ref. [130], $(\mathring{\psi}^{A},\mathring{\pi}_{A})$ in Ref. [113], and $(\Phi_{A},\alpha_{A})$ in Refs. [96, 49]. Different alphabets for the indices are also used. which is chosen to enforce the requirements that (i) the forces $f^{\alpha}_{(n)}(x^{A}_{p},u_{A})$ are $2\pi$-periodic in $\varphi_{A}$, and (ii) the evolution equations for $\varphi_{A}$ and $\pi_{A}$, obtained from the equation of motion (3), take the form
$\frac{d\varphi_{A}}{dt}$ $:=\Omega_{A}(\pi_{B},\varepsilon)=\Omega_{A}^{(0)}(\pi_{B})+\varepsilon\Omega_{A}^{(1)}(\pi_{B})+\ldots,$ (4)
$\frac{d\pi_{A}}{dt}$ $\hphantom{:}=\varepsilon\!\left[F_{A}^{(0)}(\pi_{B})+\varepsilon F_{A}^{(1)}(\pi_{B})+\ldots\right]\!.$ (5)
The motion is approximately biperiodic, with independent radial and azimuthal frequencies $\Omega_{A}=(\Omega_{r},\Omega_{\phi})$ that slowly evolve due to dissipation. The fast variables $\varphi_{A}=(\varphi_{r},\varphi_{\phi})$ describe the phases of this biperiodic motion; if dissipation were artificially ignored, the phase $\varphi_{A}$ would evolve by exactly $2\pi$ on each radial or azimuthal period $2\pi/\Omega_{A}={\cal O}(m_{1})$. The slow variables $\pi_{A}$ evolve due to dissipation on the long time scale of order $m_{1}/\varepsilon$ (and the frequencies hence evolve along with them). As we discuss in the body of this paper, there is considerable freedom in the choice of variables $\pi_{A}$, but they are typically chosen to be quasi-Keplerian parameters: a relativistic semi-latus rectum and eccentricity, $\pi_{A}=(p,e)$, which provide the simplest parameterization of the original coordinates, $x^{A}_{p}(\varphi_{B},\pi_{B},\varepsilon)$ and $p_{A}(\varphi_{B},\pi_{B},\varepsilon)$. Explicit expressions for $\Omega^{(0)}_{A}(p,e)$ are given in Eqs. (35)–(36) below.
The transformation $(x^{A}_{p},p_{A})\mapsto(\varphi_{A},\pi_{A})$ is given in Refs. [130, 113], along with expressions for $\Omega^{(1)}_{A}$ and $F^{(n)}_{A}$ in terms of the forces $f^{\alpha}_{(n)}$. In practice, such transformations are typically found starting from coordinates $(\varphi^{(0)}_{A},\pi^{(0)}_{A})$ adapted to the zeroth-order, background-geodesic case, satisfying
$\frac{d\varphi_{A}^{(0)}}{dt}=\Omega^{(0)}_{A}(\pi^{(0)}_{B})\quad\text{and}\quad\frac{d\pi^{(0)}_{A}}{dt}=0$ (6)
when $\varepsilon=0$. $\pi^{(0)}_{A}$ are constants of motion for a Schwarzschild geodesic, and $\varphi_{A}^{(0)}$ grow exactly linearly in time in that case. At finite $\varepsilon$, $d\varphi^{(0)}_{A}/dt$ and $d\pi^{(0)}_{A}/dt$ have oscillatory dependence on $\varphi^{(0)}_{A}$. The variables $(\varphi_{A},\pi_{A})$ are then found by performing a near-identity averaging transformation [158, 130, 111, 74, 113, 107],
$\varphi_{A}$ $=\varphi_{A}^{(0)}+\varepsilon\varphi_{A}^{(1)}(\varphi_{B}^{(0)},\pi^{(0)}_{B})+{\cal O}(\varepsilon^{2}),$ (7)
$\pi_{A}$ $=\pi_{A}^{(0)}+\varepsilon\pi_{A}^{(1)}(\varphi_{B}^{(0)},\pi^{(0)}_{B})+{\cal O}(\varepsilon^{2}),$ (8)
designed to ensure there are no oscillations in $d\varphi_{A}/dt$ and $d\pi_{A}/dt$.
The critical step in the multiscale expansion is to utilize these phase-space coordinates directly within the Einstein equations. We first adopt a time coordinate $s=t-\kappa(r)$ that reduces to advanced time $v=t+r^{*}$ at the horizon, Schwarzschild time $t$ at the particle, and retarded time $u=t-r^{*}$ at infinity. We then extend $(\varphi_{A},\pi_{A})$ off the worldline by taking them to be constant on slices of constant $s$ and write the small-mass-ratio expansion (1) as
$g_{\alpha\beta}=g^{(0)}_{\alpha\beta}(x^{i})+\varepsilon h^{(1)}_{\alpha\beta}(\varphi_{A},\pi_{A},x^{i})\\ +\varepsilon^{2}h^{(2)}_{\alpha\beta}(\varphi_{A},\pi_{A},x^{i})+{\cal O}(\varepsilon^{3}),$ (9)
where $x^{i}=(r,\theta,\phi)$ are the spatial coordinates on each $s=\text{constant}$ slice. Here all time dependence in the metric is encoded in its dependence on $(\varphi_{A},\pi_{A})$, making the metric a function on phase space (crossed with space); this form of the metric is derived from first principles in Ref. [107]. Taking the $r\to\infty$ limit of Eq. (9) then yields the waveform at infinity as a function of $(\varphi_{A},\pi_{A})$ and sky angles $(\theta,\phi)$. Typically the functions in Eq. (9) and in the corresponding waveform are computed and stored as coefficients in a discrete Fourier series,
$h^{(n)}_{\alpha\beta}=\sum_{k_{A}\in\mathbb{Z}^{2}}h^{(n,k_{A})}_{\alpha\beta}(\pi_{B},x^{i})e^{-ik_{A}\varphi_{A}}.$ (10)
The waveform as a function of time is then generated by solving Eqs. (4)–(5) and summing the Fourier modes [96]; [^4] In addition to Eqs. (4)–(5), 1PA accuracy also requires evolution equations for the primary’s mass and spin, but these will not be important for the present paper. For an initially nonspinning primary, the spin parameter remains small (of order $m_{2}$) throughout the inspiral [117], and it can be accounted for analytically as in Ref. [114]. this is a rapid process by virtue of the fact that the right-hand sides of Eqs. (4)–(5) change slowly, allowing large time steps as no oscillations need to be numerically resolved [158].
In this section, we have followed the traditional order counting in SF theory. An $n$th post-adiabatic (nnPA) term contributes to the phase evolution at order $\varepsilon^{n}$ relative to the leading, adiabatic (0PA) evolution; the numeric labels in Eqs. (4) and (5) represent this post-adiabatic counting. On the other hand, an $n$th-order self-force term (nnSF) is computed from metric perturbations that are of absolute order $\varepsilon^{n}$; the numeric labels in Eqs. (1)–(3) follow this counting. The two orderings differ because of the secular effect of dissipative forces: $\Omega^{(n)}_{A}$ is both nnPA and nnSF, while the dissipative forcing functions $F^{(n)}_{A}$ are nnPA but (n+1)(n+1)SF.
II.2 Hamiltonian framework and conserved quantities
The existence of the variables $(\varphi_{A},\pi_{A})$ is dependent on the fact that the zeroth-order, background-geodesic motion is an integrable Hamiltonian system [146, 86], with two independent constants of motion. A Hamiltonian analysis shows how those constants of motion naturally extend to 1SF order.
For a geodesic in Schwarzschild spacetime, the spacetime’s Killing symmetries imply the energy
$E_{(0)}(p_{A})=-p_{t}$ (11)
and angular momentum
$L_{(0)}(p_{A})=p_{\phi}$ (12)
are constants. These constants of motion, $P^{(0)}_{A}=(E_{(0)},L_{(0)})$, are in one-to-one correspondence with the quasi-Keplerian parameters $\pi_{A}=(p,e)$; the explicit geodesic relationships $P^{(0)}_{A}(\pi_{B})$ are given in Eq. (30) below. While these constants follow from spacetime symmetries, we can also obtain them from Hamiltonian methods: the Schwarzschild geodesic equation can be written as Hamilton’s equations on the 4D phase space [107],
$\frac{dx^{A}_{p}}{dt}$ $=\frac{\partial H^{(0)}_{\rm 4D}}{\partial p_{A}},$ (13)
$\frac{dp_{A}}{dt}$ $=-\frac{\partial H^{(0)}_{\rm 4D}}{\partial x^{A}_{p}},$ (14)
with the Hamiltonian
$H^{(0)}_{\rm 4D}(x^{A}_{p},p_{A})=\sqrt{-\frac{g^{AB}_{(0)}(r_{p})p_{A}p_{B}+(m_{2})^{2}}{g^{tt}_{(0)}(r_{p})}}.$ (15)
Since the Hamiltonian is independent of $t$, its on-shell value is constant, equal to the energy (11), as one can see by rearranging $g^{\alpha\beta}_{(0)}p_{\alpha}p_{\beta}=-(m_{2})^{2}$. Since the Hamiltonian is independent of $\phi_{p}$, the angular momentum (12) is also constant: $dp_{\phi}/dt=-\partial H^{(0)}_{\rm 4D}/\partial\phi_{p}=0$.
To reach the variables $\varphi^{(0)}_{A}$, we note that for bound, bi-periodic geodesic orbits, surfaces of constant $P^{(0)}_{A}$ define 2D tori in phase space, and the bi-periodic orbits wrap around the surfaces of these tori. We can define action variables $J^{(0)}_{A}=(J_{r(0)},J_{\phi(0)})$ as integrals over the tori,
$J^{(0)}_{A}=\frac{1}{2\pi}\oint_{{\cal C}^{(0)}_{A}}p_{B}\,dx^{B}_{p},$ (16)
where ${\cal C}^{(0)}_{A}$ denotes one of the two topologically distinct loops on the torus. The action variables $J^{(0)}_{A}$ are canonically conjugate (at 0SF) to the geodesic action-angles $\varphi^{(0)}_{A}$, meaning the geodesic frequencies can be written as $d\varphi_{A}^{(0)}/dt:=\Omega^{(0)}_{A}=\partial H^{(0)}_{\rm 4D}/\partial J^{(0)}_{A}$. Since $p_{\phi}=L^{(0)}$ is constant on each torus, we also have $J^{(0)}_{\phi}=L^{(0)}$. $J_{r(0)}$ can also be written as a function of $P^{(0)}_{A}$; see Eq. (32) below. One can then work with whichever set of constants is convenient: $\pi^{(0)}_{A}$, $P^{(0)}_{A}$, or $J^{(0)}_{A}$.
Reference [107] extended this analysis to the full 1PA dynamics, including dissipation. Specifically, Ref. [107] showed that the evolution equations (4)–(5) are equivalent (at 1PA order) to
$\frac{d\varphi_{A}}{dt}$ $=\frac{\partial H_{\rm 4D}}{\partial J_{A}},$ (17)
$\frac{dJ_{A}}{dt}$ $=\varepsilon\left[G^{(0)}_{A}(J_{B})+\varepsilon G^{(1)}_{A}(J_{B})\right],$ (18)
where
$H_{\rm 4D}(J_{A},\varepsilon)=E_{(0)}(J_{A})-\frac{1}{4}\varepsilon m_{2}\left\langle\frac{d\tau}{dt}h_{uu}^{\mathrm{R}(1)}\right\rangle(J_{A})$ (19)
is a Hamiltonian that governs the 1SF conservative sector, and $G^{(n)}_{A}$ are dissipative forcing functions that can be written in terms of a pseudo-Hamiltonian (a certain two-point function on phase space that will not be needed in this paper). Here we have introduced the shorthand $h_{uu}^{\mathrm{R}(1)}:=h_{\alpha\beta}^{\mathrm{R}(1)}u^{\alpha}u^{\beta}$ as well as angular brackets to denote a torus average: $\langle\cdot\rangle:=\frac{1}{(2\pi)^{2}}\oint\cdot\,d^{2}\varphi$. Due to the axial symmetry of the background, all functions we average will be independent of $\varphi_{\phi}$ when evaluated on the worldline, reducing the average to an average over one radial cycle, $\langle\cdot\rangle=\frac{1}{2\pi}\oint\cdot d\varphi_{r}$. We can equivalently write this as a time average over a radial period $T_{r}$: for any function $f(\varphi_{r},J_{A})$, we have
$\langle f\rangle(J_{A})=\frac{1}{T_{r}}\int_{0}^{T_{r}}f(\Omega_{r}t,J_{A})dt.$ (20)
The action variables can equivalently be replaced by quasi-Keplerian parameters $\pi_{A}$ in this equality.
The action variables $J_{A}$ are the unique variables canonically conjugate to $\varphi^{A}$ with respect to the Hamiltonian $H_{\rm 4D}$. In the conservative sector they are constants, implying that conservative 1SF motion is restricted to tori of constant $J_{A}$, wtih $\varphi_{A}$ as coordinates on each torus. $J_{A}$ can trivially be written as integrals over these tori:
$J_{A}=\frac{1}{2\pi}\oint_{{\cal C}_{A}}J_{B}d\varphi_{B}=\frac{1}{2\pi}\oint_{{\cal C}_{A}}p^{\rm can}_{B}dx^{B}_{p},$ (21)
where ${\cal C}_{A}$ are the two independent loops on the torus, and $p^{\rm can}_{B}$ (which we will not require) is canonically conjugate to $x^{A}_{p}$ with respect to $H_{\rm 4D}$. In the dissipative system, the systems slowly evolves from one to torus to the next.
This Hamiltonian construction provides natural definitions of energy and angular momentum. As at 0SF order, the 1SF energy $E$ is the on-shell value of $H_{\rm 4D}$ (which, note, is simply equal to $H_{\rm 4D}$’s value as a function of the action variables, since the Hamiltonian is independent of the angle variables). Similarly, the angular momentum is the azimuthal action variable, $L=J_{\phi}$. We can again work with any choice of variables $\pi_{A}$, $P_{A}=(E,L)$, or $J_{A}$.
The conservative dynamics is encoded in the torus-averaged Detweiler-Barack-Sago redshift [^5] We avoid labeling the terms here with $(0)$ and $(1)$ because they do not correspond to an expansion in powers of $\varepsilon$ at fixed phase-space coordinates. $z$ and $z_{\rm Schw}$ represent the two redshifts $d\tilde{\tau}/dt$ and $d\tau/dt$ along the same accelerated orbit. $\langle z\rangle=\langle d\tilde{\tau}/dt\rangle=\langle z_{\rm Schw}+\varepsilon z_{\rm 1SF}\rangle$ [65, 19], where $\tilde{\tau}$ is proper time in the effective metric $\tilde{g}_{\alpha\beta}$, $z_{\rm Schw}=d\tau/dt=1/u^{t}$, and
$z_{\rm 1SF}=-\frac{1}{2}z_{\rm Schw}h_{uu}^{\mathrm{R}(1)}.$ (22)
We can use the averaged redshift as a generating function for all of our quantities of interest. For this purpose it is useful to expand $\langle z\rangle$ in powers of $\varepsilon$ at fixed values of the frequencies, where we use $(\Omega_{A},\varphi_{A})$ as phase-space coordinates and write $x^{A}_{p}=x^{A}_{(0)}(\Omega_{B},\varphi_{B})+\varepsilon x^{A}_{(1)}(\Omega_{B},\varphi_{B})$, leading to
$\langle{z}\rangle({\Omega}_{A},\varepsilon)=\langle{z}_{(0)}\rangle({\Omega}_{A})+\varepsilon\langle{z}_{(1)}\rangle({\Omega}_{A}).$ (23)
Here $x^{A}_{(0)}$ is the same function of $(\Omega_{A},\varphi_{A})$ as for a geodesic with those frequencies, $\langle{z}_{(0)}\rangle$ is $\langle z_{\rm Schw}\rangle$ evaluated at $x^{A}_{p}=x^{A}_{(0)}$, and
$\langle z_{(1)}\rangle=-\frac{1}{2}\langle z_{\rm Schw}h_{uu}^{\mathrm{R}(1)}\rangle.$ (24)
The principle advantage of this expansion at fixed frequencies is that there are no additional terms at linear order in $\varepsilon$, while expansions at fixed values of other phase-space coordinates will generically contain additional ${\cal O}(\varepsilon)$ terms from the expansion of $z_{\rm Schw}$; see Appendix D of Ref. [107].
It will be useful to work with dimensionless functions of dimensionless arguments rather than leaving functional dependence on $m_{1}$ implicit, as we have done up to this point. We define reduced, dimensionless counterparts to energy $E$, $L$, and $J_{r}$ via
$E$ $=m_{2}\hat{E}$, $L$ $=m_{1}m_{2}\hat{L}$, $J_{r}$ $=m_{1}m_{2}\hat{J}_{r},$ (25)
along with dimensionless frequencies $\hat{\Omega}_{A}=m_{1}\Omega_{A}$. Expanding $\hat{P}_{A}$ and $\hat{J}_{A}$ in powers of $\varepsilon$ at fixed $\hat{\Omega}_{A}$ yields
$\hat{E}(\hat{\Omega}_{A},\varepsilon)$ $=\hat{E}_{(0)}(\hat{\Omega}_{A})+\varepsilon\hat{E}_{(1)}(\hat{\Omega}_{A}),$ (26a)
$\hat{L}(\hat{\Omega}_{A},\varepsilon)$ $=\hat{L}_{(0)}(\hat{\Omega}_{A})+\varepsilon\hat{L}_{(1)}(\hat{\Omega}_{A}),$ (26b)
$\hat{J}_{r}(\hat{\Omega}_{A},\varepsilon)$ $=\hat{J}_{r(0)}(\hat{\Omega}_{A})+\varepsilon\hat{J}_{r(1)}(\hat{\Omega}_{A}),$ (26c)
where there is now no hidden dependence on $m_{1}$. The 1SF terms in these expansions, which will be the focus of this paper, are given in terms of the redshift by [107]
$\hat{E}_{(1)}(\hat{\Omega}_{A})$ $=\frac{1}{2}\left(\langle z_{(1)}\rangle-\hat{\Omega}_{A}\,\frac{\partial\langle z_{(1)}\rangle}{\partial\hat{\Omega}_{A}}\right),$ (27a)
$\hat{L}_{(1)}(\hat{\Omega}_{A})$ $=-\frac{1}{2}\frac{\partial\langle z_{(1)}\rangle}{\partial\hat{\Omega}_{\phi}},$ (27b)
$\hat{J}_{r(1)}(\hat{\Omega}_{A})$ $=-\frac{1}{2}\frac{\partial\langle z_{(1)}\rangle}{\partial\hat{\Omega}_{r}},$ (27c)
where $\langle z_{(1)}\rangle=\langle z_{(1)}\rangle(\hat{\Omega}_{A})$.
III Geodesic order
In this section we review the case of geodesic orbits in more detail, particularly the quasi-Keplerian parameterization and the relationships between the various constants of motion, largely following Ref. [18]. We also carry out PN expansions of the geodesic relationships, defining alternative frequency variables that are useful for that purpose.
III.1 Exact expressions
Rather than writing the geodesic motion directly in terms of the phase-space coordinates $(\varphi_{A},\pi_{A})$ used in waveform generation, we can more straightforwardly write it in terms of the (Darwin) relativistic anomaly $\chi$. In terms of this radial phase, Darwin’s quasi-Keplerian parameterization of the geodesic radial motion is [62]
$r_{(0)}(\chi,\pi_{A})$ $=\frac{pm_{1}}{1+e\cos\chi}\,.$ (28)
The orbital parameters $(p,e)$ are expressible in terms of the (geodesic) radii of periastron $r_{\rm min}$ and apastron $r_{\rm max}$ as
$p$ $=\frac{2r_{\rm min}r_{\rm max}}{m_{1}(r_{\rm max}+r_{\rm min})}\,$, $e$ $=\frac{r_{\rm max}-r_{\rm min}}{r_{\rm max}+r_{\rm min}}\,.$ (29)
The constants of motion can then be expressed in terms of $p$ and $e$. Explicitly, the reduced energy and angular momentum are
$\hat{E}_{(0)}(\pi_{A})$ $=\sqrt{\frac{(p-2+2e)(p-2-2e)}{p(p-3-e^{2})}}\,$, (30a)
$\hat{L}_{(0)}(\pi_{A})$ $=\frac{p}{\sqrt{p-3-e^{2}}}\,$, (30b)
and we define the semilatus rectum at the separatrix as
$p_{\star}=6+2e\,.$ (31)
In the region of parameter space where $p>p_{\star}$ and $0<e<1$, the determinant of the Jacobian matrix between $(\hat{E}_{(0)},\hat{L}_{(0)})$ and $(p,e)$ is regular, so the map between these two pairs of variables is one-to-one. The reduced radial action $\hat{J}_{r(0)}(\pi_{A})$ is given by Eq. (16):
$\hat{J}_{r(0)}=\frac{1}{\pi}\int_{\hat{r}_{\rm min}}^{\hat{r}_{\rm max}}d\hat{r}\,\frac{\sqrt{\hat{r}^{4}\hat{E}^{2}-(\hat{r}^{2}-2\hat{r})(\hat{L}^{2}+\hat{r}^{2})}}{\hat{r}^{2}-2\hat{r}}\,$, (32)
where we defined the dimensionless radius $\hat{r}=r_{(0)}/m_{1}$ and used $g^{rr}_{(0)}u_{r}^{2}=-1-g^{tt}_{(0)}u_{t}^{2}-g^{\phi\phi}_{(0)}u_{\phi}^{2}$ to rewrite the integrand.
The evolution of the coordinate time, azimuthal phase, and proper time along the orbit can be written as
$\frac{\mathrm{d}t_{(0)}}{\mathrm{d}\chi}$ $=\frac{m_{1}\,p^{2}}{(p-2-2e\cos\chi)(1+e\cos\chi)^{2}}$
$\qquad\times\sqrt{\frac{(p-2-2e)(p-2+2e)}{p-6-2e\cos\chi}}\,$, (33a)
$\frac{\mathrm{d}\phi_{(0)}}{\mathrm{d}\chi}$ $=\sqrt{\frac{p}{p-6-2e\cos\chi}}\,$, (33b)
$\frac{\mathrm{d}\tau_{(0)}}{\mathrm{d}\chi}$ $=\frac{m_{1}\,p^{3/2}}{(1+e\cos\chi)^{2}}\sqrt{\frac{p-3-e^{2}}{p-6-2e\cos\chi}}\,.$ (33c)
Thus, the cumulative change of these times and phases over one radial period are
$T_{r(0)}$ $=\!\int_{0}^{2\pi}\!\!\mathrm{d}\chi\,\frac{\mathrm{d}t_{(0)}}{\mathrm{d}\chi}\,$, (34a)
$\Phi_{(0)}$ $=\!\int_{0}^{2\pi}\!\!\mathrm{d}\chi\,\frac{\mathrm{d}\phi_{(0)}}{\mathrm{d}\chi}\,$, (34b)
$\mathcal{T}_{r(0)}$ $=\!\int_{0}^{2\pi}\!\!\mathrm{d}\chi\,\frac{\mathrm{d}\tau_{(0)}}{\mathrm{d}\chi}\,.$ (34c)
These definite integrals can be expressed in terms of elliptic integrals [75, 158]. Introducing the reduced periods $\hat{T}_{r}=T_{r}/m_{1}$ and $\hat{\mathcal{T}}_{r}=\mathcal{T}_{r}/m_{1}$ and restricting to the case of bound stable orbits $p>p_{\star}$, we find that at geodesic order they are given by
$\hat{T}_{r(0)}$ $=\frac{2p\,a_{+}\,a_{-}\,b_{+}}{(1-e^{2})(p-4)}\mathrm{E}(m_{r})-\frac{2pa_{+}a_{-}}{(1-e^{2})b_{+}}\mathrm{K}(m_{r})$
$+\frac{4a_{+}\,a_{-}\!\left[p(p-1-3e^{2})\!-\!8(1-e^{2})\right]}{(1-e)(1-e^{2})(p-4)b_{+}}\Pi\!\left(\!-\frac{2e}{1-e}\Big|m_{r}\!\!\right)$
$+\frac{16a_{-}}{a_{+}b_{+}}\Pi\left(\frac{4e}{p-2+2e}\Big|m_{r}\right),$ (35a)
$\Phi_{(0)}$ $=\frac{4\sqrt{p}}{b_{+}}\mathrm{K}(m_{r}),$ (35b)
$\hat{\mathcal{T}}_{r(0)}$ $=\frac{2p^{3/2}b_{+}c_{-}}{(1-e^{2})(p-4)}\mathrm{E}\left(m_{r}\right)-\frac{2p^{3/2}c_{-}}{(1-e^{2})b_{+}}\mathrm{K}\left(m_{r}\right)$
$+\frac{4p^{3/2}c_{-}^{3}}{(1-e^{2})(1-e)(p-4)b_{+}}\Pi\left(-\frac{2e}{1-e}\Big|m_{r}\right),$ (35c)
where we have introduced the shorthands $a_{\pm}=\sqrt{p-2\pm 2e}$, $b_{\pm}=\sqrt{p-6\pm 2e}$, $c_{\pm}=\sqrt{p-3\pm e^{2}}$, and $m_{r}=4e/b_{+}^{2}$. The elliptic integrals $\mathrm{E}(\cdot)$, $\mathrm{K}(\cdot)$, and $\Pi(\cdot|\cdot)$ are defined in Appendix A for completeness. To our knowledge, the above expression for $\mathcal{T}_{r}$ has never appeared in the literature, but the expressions for $T_{r}$ and $\Phi$ were known [158].
From the coordinate-time and proper-time periods and elapsed azimuthal angle, we can also define the orbital frequencies and orbit-averaged redshift
$\hat{\Omega}_{r}$ $=\frac{2\pi}{\hat{T}_{r}},$ $\hat{\Omega}_{\phi}$ $=\frac{\Phi}{\hat{T}_{r}},$ $\langle z_{\rm Schw}\rangle$ $=\frac{\hat{\mathcal{T}}_{r}}{\hat{T}_{r}}.$ (36)
At geodesic order these are expressed in terms of $(p,e)$ using the formulas for $\hat{T}_{r(0)}$, $\Phi_{(0)}$, and $\hat{\cal T}_{r(0)}$ above:
$\hat{\Omega}_{r}^{(0)}$ $=\frac{2\pi}{\hat{T}_{r}^{(0)}},$ $\hat{\Omega}_{\phi}^{(0)}$ $=\frac{\Phi^{(0)}}{\hat{T}_{r}^{(0)}},$ $\langle z_{(0)}\rangle$ $=\frac{\hat{\mathcal{T}}_{r}^{(0)}}{\hat{T}_{r}^{(0)}}.$ (37)
Note that alternative formulas for the frequencies exist in the literature (e.g., in the Black Hole Perturbation Toolkit [34]); they can be reconciled with our expressions using the relations (114)–(115) between elliptic integrals.
The reduced radial action (32) can also be written directly in terms of elliptic integrals, as in Ref. [170]. Equivalently, we can write it in terms of quantities given above by using the normalization condition $\hat{\omega}^{(0)}_{\alpha}\hat{J}^{(0)}_{\alpha}=-1$ [76], where $\hat{\omega}^{(0)}_{\alpha}:=\langle z_{\rm(0)}\rangle^{-1}\hat{\Omega}^{(0)}_{\alpha}$ are the reduced orbital frequencies with respect to proper time, $\hat{J}_{t(0)}:=-\hat{E}_{(0)}$, and $\hat{\Omega}^{(0)}_{t}:=1$. Rearranging the normalization condition yields
$\hat{J}_{r(0)}=\frac{1}{2\pi}\left(\hat{T}_{r(0)}\ \hat{E}_{(0)}-\Phi_{(0)}\hat{L}_{(0)}-\hat{\mathcal{T}}_{r(0)}\right)\,.$ (38)
We have checked that this expression is in agreement with Eq. (34) of Ref. [170] in the appropriate limit.
The parameters $(p,e)$ as defined in (29) provide globally valid phase-space coordinates everywhere outside the separatrix between bound and plunging orbits, which lies on the curve $p=p_{\star}$. For given initial values of $\phi_{(0)}$ and $\chi_{(0)}$, $p$ and $e$ uniquely specify a geodesic. However, even when restricting to stable bound orbits, for given initial $\phi_{(0)}$ and $\chi_{(0)}$, there exist two physically inequivalent orbits associated with the pair of frequencies $(\Omega_{r},\Omega_{\phi})$; this is the so-called isofrequency pairing [163]. The isofrequency degeneracy curve, which separates the parameter space into two domains of such inequivalent orbits, is given by the condition
$\mathrm{det}\,\bm{J}=0\,$, (39)
where the Jacobian matrix is defined by
$\bm{J}=\begin{pmatrix}\frac{\partial\hat{\Omega}_{r}}{\partial p}&\frac{\partial\hat{\Omega}_{r}}{\partial e}\\ \frac{\partial\hat{\Omega}_{\phi}}{\partial p}&\frac{\partial\hat{\Omega}_{\phi}}{\partial e}\end{pmatrix}\,$. (40)
This singular curve in parameter space will be highly relevant in our calculations of 1SF quantities. We will denote the relation between $p$ and $e$ on the isofrequency curve by $p_{\rm iso}(e)$.
III.2 Post-Newtonian expansion
The PN expansion is an expansion in small velocity, or equivalently, large separation. Here, it will translate to an expansion for $p\gg 1$ at fixed $e$ in the range $0\leq e<1$. Alternatively, we can introduce another set of parameters that are more directly related to the frequencies:
$y=\hat{\Omega}_{\phi}^{2/3}$ and $\lambda=\frac{3y}{\hat{\Omega}_{\phi}/\hat{\Omega}_{r}-1}\,$, (41)
such that the PN expansion now corresponds to $y\ll 1$ with $0<\lambda\leq 1$.
Starting from (30) and (36), we perform a PN expansion of the fundamental frequencies, the redshift, and the constants of motion. For this, we use the expansions of the elliptic integrals provided in Eq. (113). The first few PN orders read
$\hat{\Omega}^{(0)}_{r}$ $=\left(\frac{1-e^{2}}{p}\right)^{3/2}\left[1-\frac{3(1-e^{2})}{p}+\mathcal{O}\!\left(\frac{1}{p^{2}}\right)\right],$ (42a)
$\hat{\Omega}^{(0)}_{\phi}$ $=\left(\frac{1-e^{2}}{p}\right)^{3/2}\left[1+\frac{3e^{2}}{p}+\mathcal{O}\!\left(\frac{1}{p^{2}}\right)\right],$ (42b)
$\langle z_{(0)}\rangle$ $=1-\frac{3(1-e^{2})}{2p}+\mathcal{O}\!\left(\frac{1}{p^{2}}\right)\,$, (42c)
$\hat{E}_{(0)}$ $=-\frac{1-e^{2}}{2p}+\frac{3(1-e^{2})^{2}}{8p^{2}}+\mathcal{O}\!\left(\frac{1}{p^{3}}\right)\,$, (42d)
$\hat{L}_{(0)}$ $=\sqrt{p}+\frac{3+e^{2}}{2\sqrt{p}}+\mathcal{O}\!\left(\frac{1}{p^{3/2}}\right)\,$, (42e)
$\hat{J}_{r(0)}$ $=\sqrt{p}\left[\frac{1}{\sqrt{1-e^{2}}}-1\right]+\mathcal{O}\!\left(\frac{1}{\sqrt{p}}\right)\,.$ (42f)
The expansions are provided in the ancillary file PN_expressions.csv [1] at relative 12PN for $\hat{\Omega}_{r}^{(0)}$ and $\hat{\Omega}_{\phi}^{(0)}$, 10PN for $\langle z_{(0)}\rangle$ and 9PN for $\hat{E}_{(0)}$, $\hat{L}_{(0)}$, and $\hat{J}_{r(0)}$. This translates to (geodesic-order) expansions for the reduced frequency variables which read
$y_{(0)}$ $=\frac{1-e^{2}}{p}\left[1+\frac{2e^{2}}{p}+\mathcal{O}\!\left(\frac{1}{p^{2}}\right)\right],$ (43a)
$\lambda_{(0)}$ $=(1-e^{2})\left[1+\frac{1}{p}\left(-\frac{9}{2}+\frac{7}{4}e^{2}\right)+\mathcal{O}\!\left(\frac{1}{p^{2}}\right)\right].$ (43b)
The expansions for $y$ and $\lambda$ are provided, respectively, at 12PN and 11PN order in the ancillary file PN_expressions.wl [1]. Indeed, there is a loss of one PN order in $\lambda$ due to the degeneracy between radial and azimuthal frequencies at Newtonian order; see [152] for a discussion. Inverting the latter expansion, one finds that
[…]
where these expansion for $p$ and $e$ are provided, respectively, at 12PN and 11PN order in the in the ancillary file PN_expressions.wl [1].
Our 1SF results will focus on parameterizations in terms of $(p,e)$, and our comparisons with numerical results will indicate that the PN expansion is significantly more accurate when written as a large-$p$ expansion at fixed $e$ rather than a small-$y$ expansion at fixed $\lambda$. However, comparisons with results in strict PN theory are far simpler in terms of $(y,\lambda)$. For this reason, we also provide high-order PN expansions in terms of $(y,\lambda)$ in the ancillary file PN_expressions.wl [1].
IV First order
In this section we numerically and analytically compute the 1SF corrections to the conserved quantities $(\hat{E},\hat{L},\hat{J}_{r})$ as functions of the frequencies $\hat{\Omega}_{A}$ across much of the bound-orbit parameter space, starting from the formulas (27) that express these functions in terms of the 1SF redshift correction $\langle z_{(1)}\rangle$. We obtain our numerical results on a dense grid up to $e=0.6$ and $p=p_{\star}+200$, suitable for use in waveform-generation codes, and perform a detailed comparison with the analytical, 9PN results across that range. Our 9PN results also establish consistency with the 4PN results obtained independently within PN theory in Ref. [152].
IV.1 Quasi-Keplerian variables at 1SF order
We are ultimately interested in the expressions (or numerical values) for the constants of motion as functions of the orbital frequencies, which represent gauge-invariant relationships between gauge-invariant quantities. But it is still useful to make use of the intermediate orbital parameters $\pi_{A}=(p,e)$. These quantities provide an intuitive geometric description of the leading-order dynamics, and, therefore, are a natural choice for tiling the parameter space and for parametrizing 1SF calculations. (See Sec. IV.2.) They are also the parameters used to store offline data for practical waveform generation, as reviewed in Sec. II. Finally, as we will see in Sec. IV.2, the PN expansion of the redshift is numerically more accurate when using $(p,e)$ rather than other parameters such as $(y,\lambda)$.
However, there is considerable freedom in specifying the meaning of $(p,e)$ at 1SF order. Specifically, the multiscale expansion and waveform-generation framework are invariant under a transformation $\pi_{A}\to\pi_{A}+\varepsilon\delta\pi_{A}(\hat{J}_{B})$ (i.e., a near-identity transformation that is independent of the angles $\varphi_{A}$[113]. This freedom was explored in the context of multiscale expansions of the full, dissipative system in Refs. [158, 130, 113, 107], for example. In Refs. [127, 113], two choices were highlighted that are particularly natural for us: [^6] Another choice is to define $(p,e)$ in terms of the radial turning points of the perturbed dynamics. This is a natural extension of their geodesic definitions, but it depends on the choice of background coordinates, is only well defined in the absence of dissipation, and ties $(p,e)$ to the gauge of the metric perturbation $h^{(1)}_{\mu\nu}$, making it difficult for cross-gauge comparisons or determining their relations to $(\hat{E},\hat{L},\hat{J}_{r})$ or $(\hat{\Omega}_{r},\hat{\Omega}_{\phi})$.
  1. 1.
[…]
  2. 2.
    the fixed frequencies gauge, in which $(p,e)$ are related to $(\hat{\Omega}_{r},\hat{\Omega}_{\phi})$ by the Schwarzschild-geodesic relationships (36).
In other words, in the fixed-constants gauge, we have $\hat{P}_{A}(\pi_{B},\varepsilon)=\hat{P}^{(0)}_{A}(\pi_{B})$, independent of $\varepsilon$; in the fixed-frequencies gauge, $\hat{\Omega}_{A}(\pi_{B},\varepsilon)=\hat{\Omega}^{(0)}_{A}(\pi_{B})$.
In this work we choose the latter, making it straightforward to parametrize $(\hat{E},\hat{L},\hat{J}_{r})$ with $(p,e)$ in place of $(\hat{\Omega}_{r},\hat{\Omega}_{\phi})$ in Eq. (26); this definition was previously adopted, e.g., in Eq. (71) of Ref. [19]. The fixed-constants gauge has the analogous advantage when calculating $\hat{\Omega}_{A}$ as a function of $\hat{P}_{A}$, rather than the converse:
$\hat{\Omega}_{A}(\hat{P}_{B},\varepsilon)=\hat{\Omega}^{(0)}_{A}(\hat{P}_{B})+\varepsilon\,\hat{\Omega}^{(1)}_{A}(\hat{P}_{B}).$ (45)
Here we can use $(p,e)$ in place of $\hat{P}_{A}$ if we adopt a fixed-constant gauge.
We focus on calculations of the corrections to the constants of motion at fixed frequencies, $\hat{P}^{(1)}_{A}(\hat{\Omega}_{B})$, but for waveform generation the converse, $\hat{\Omega}^{(1)}_{A}(\hat{P}_{B})$, might be more useful; see Sec. V.4 below. Relating the two is straightforward. Substituting Eq. (45) in Eq. (26) and expanding, we obtain
$\Omega^{(1)}_{A}(\hat{P}_{B})=-\frac{\partial\hat{\Omega}^{(0)}_{A}}{\partial\hat{P}_{B}}\hat{P}^{(1)}_{B}$. (46)
IV.2 Redshift invariant
We compute $h_{\mu\nu}^{\mathrm{R}(1)}$, and subsequently $\langle z_{(1)}\rangle$, in a Fourier decomposition and an expansion in spherical harmonics, using the Python library pybhpt [^7] https://pybhpt.readthedocs.io/en/latest/ [122] and the methods outlined in Ref. [121]. Crucially, with this approach, $\langle z_{(1)}\rangle$ reduces to a formally infinite sum over the spherical-harmonic $l$-modes,
$\langle z_{(1)}\rangle=\sum_{\ell=0}^{\infty}\langle z^{\ell}_{(1)}\rangle.$ (47)
The convergence of the sum is impacted by the regularization procedure that is employed to obtain $h_{uu}^{\mathrm{R(1),\ell}}$ and subsequently $\langle z^{\ell}_{(1)}\rangle$. For the mode-sum regularization used in Ref. [121], $\langle z^{\ell}_{(1)}\rangle\sim\ell^{-2}$ as $\ell\rightarrow\infty$. Thus, truncating the mode sum at some finite $\ell_{\mathrm{max}}$ introduces an error $\sim\ell_{\mathrm{max}}^{-1}$. To accelerate the convergence of the mode-sum and improve the truncation error, we employ the same large-$\ell$ fitting procedure as described in Ref. [121], which produces an estimated value and error for $\langle z_{(1)}\rangle$. This fitting error is the dominant source of error in $\langle z_{(1)}\rangle$.
Consequently, we store the $\ell$-modes $\langle z^{\ell}_{(1)}\rangle$, the fitted value of $\langle z_{(1)}\rangle$, and its estimated error on a two-dimensional grid in $(p,e)$. Defining $\delta\equiv p-6-2e$, the bounds of the domain are set by
$\delta_{\mathrm{min}}$ $=10^{-3},$ $\delta_{\mathrm{max}}$ $=2\times 10^{2},$ $e_{\mathrm{max}}=0.6,$
so that $\delta\in[\delta_{\mathrm{min}},\delta_{\mathrm{max}}]$ and $e\in[0,e_{\mathrm{max}}]$. Rather than sampling directly in $(p,e)$, we make use of intermediate coordinates $(u,w)$ defined by
$u(p,e)$ $=\frac{\log\delta-\log\delta_{\mathrm{min}}}{\log\delta_{\mathrm{max}}-\log\delta_{\mathrm{min}}},$ (48a)
$w(p,e)$ $=\left[\frac{e}{e_{\mathrm{max}}}\right]^{1/2},$ (48b)
to better capture the behavior near the separatrix.
The averaged redshift invariant is also known analytically as a PN expansion. It is provided in terms of $(p,e)$ at 10PN order in [118]; see Sec. V.B. for the 8.5PN result and the Black Hole Perturbation Toolkit [34] for the full 10PN expression (we reproduce the full result in our ancillary file PN_expressions.wl [1]). The conventions in that reference are the same as here: fixed frequency, with $(p,e)$ geodesically related to $(\Omega_{r},\Omega_{\phi})$. The redshift is exact in eccentricity up to 3.5PN; at 4PN, it is expressed in terms of an enhancement function $\Lambda_{0}(e)$, which is known exactly as an infinite sum over Bessel modes and whose resummation is rigorously established and very accurate for all values of the eccentricity [118, 152]; beyond 4.5PN, some elements are controlled exactly and others are empirically resummed. The final, resummed redshift is expressed in terms of regular functions that can be safely replaced by their small-$e$ expansion without spoiling the large-$e$ behavior of the redshift itself: for instance, the chi4PNConv function corresponds to $\frac{3}{2}\lambda_{0}(e)$ as defined in [152]. In this work, we replaced all of these regular functions by their expansions up to $e^{20}$. The PN-expanded redshift finally reads
$\langle z_{(1)}\rangle=\frac{1-e^{2}}{p}\left\{1-\frac{1-e^{2}}{p}+\mathcal{O}\left(\frac{1}{p^{2}}\right)\right\},$ (49)
and we provide the full 10PN expression in the ancillary file PN_expressions.wl [1]. Using the PN map provided in Eq. 43, we re-express the redshift invariant in terms of the reduced frequencies $(y,\lambda)$. It reads
$\langle z_{(1)}\rangle=y\left\{1+y\left(1-\frac{2}{\lambda}\right)+\mathcal{O}(y^{2})\right\},$ (50)
and we provide the full 10PN expression in the ancillary file PN_expressions.wl [1]. Note that this result agrees perfectly (at linear order in the mass ratio) with the 4PN redshift recently computed by one of us using traditional PN theory [152]. We additionally provide the partial derivatives of the redshift, namely $\partial_{p}\langle z\rangle$ and $\partial_{e}\langle z\rangle$, in the ancillary file PN_expressions.wl [1].
Since the 1SF corrections to the constants of motion involve derivatives of the redshift, we also need to compute the latter numerically from our data. To compute numerical derivatives with respect to $p$ and $e$, we first subtract the 10PN expressions and normalize by the leading-order behavior, leading to
$Z_{(1)}=\frac{p}{(1-e^{2})}\left[\langle z_{(1)}\rangle-\langle z_{(1)}\rangle^{\mathrm{10PN}}\right].$ (51)
We then compute $\partial_{u}Z_{(1)}$ and $\partial_{w}Z_{(1)}$ using finite-difference stencils applied to the values of $Z_{(1)}$ on the $(u,w)$ grid. Two sources of uncertainty are considered:
  1. 1.
    propagation of fitting uncertainties in $\langle z_{(1)}\rangle$, and
  2. 2.
    finite-difference truncation errors arising from the stencil order and step size.
The fitting uncertainties are propagated in quadrature through the finite-difference stencils. Finite-difference errors are estimated by computing derivatives using four stencils: fourth-order stencils with step sizes $(\Delta u,\Delta w)$ and $(2\Delta u,2\Delta w)$, and sixth-order stencils with the same step sizes. If all four results are consistent within their uncertainties, Richardson extrapolation is applied to the two fourth-order estimates to obtain the final derivatives, with uncertainties propagated in quadrature. If the four results are not consistent, the finite-difference uncertainty is estimated from the variance among the four estimates, and Richardson extrapolation of the two sixth-order results is used to determine the final derivative values at each grid point.
After computing $\partial_{u}Z_{(1)}$ and $\partial_{w}Z_{(1)}$, it is straightforward to obtain
$\partial_{p}Z_{(1)}$ $=\frac{\partial u}{\partial p}\partial_{u}Z_{(1)},$ (52a)
$\partial_{e}Z_{(1)}$ $=\frac{\partial u}{\partial e}\partial_{u}Z_{(1)}+\frac{\partial w}{\partial e}\partial_{w}Z_{(1)},$ (52b)
and likewise,
$\partial_{p}\langle z_{(1)}\rangle$ $=\frac{(1-e^{2})}{p}\partial_{p}Z_{(1)}-\frac{(1-e^{2})}{p^{2}}Z_{(1)}+\partial_{p}\langle z_{(1)}\rangle^{\mathrm{PN}},$ (53a)
$\partial_{e}\langle z_{(1)}\rangle$ $=\frac{(1-e^{2})}{p}\partial_{e}Z_{(1)}-\frac{2e}{p}Z_{(1)}+\partial_{e}\langle z_{(1)}\rangle^{\mathrm{PN}}.$ (53b)
Once again we propagate errors by quadrature to estimate the numerical uncertainty in $\partial_{p}\langle z_{(1)}\rangle$ and $\partial_{e}\langle z_{(1)}\rangle$.
We find that derivatives with respect to $e$ require separate treatment at $e=0$. Because $\partial w/\partial e$ is singular at this point, we instead compute numerical derivatives with respect to $w^{2}$ using the same procedure as above, employing the highest-order finite-difference stencil for the final derivative values rather than Richardson extrapolation. We find $\partial_{e}\langle z_{(1)}\rangle(p,e=0)=0$ within numerical error for all sampled values of $p$, as expected.
Because $\partial_{e}\langle z_{(1)}\rangle$ vanishes at $e=0$, the corrections $\hat{E}_{(1)}$, $\hat{L}_{(1)}$, and $\hat{J}_{r(1)}$ instead depend on $\partial_{e^{2}}\langle z_{(1)}\rangle=2\,\partial^{2}_{e}\langle z_{(1)}\rangle$ in this limit; see Appendix C for further discussion. However, we find that the finite-difference scheme on the original grid becomes numerically unstable near $e=0$, in part due to a delicate cancellation of the two terms on the right-hand side of Eq. 52b as $e\to 0$. This leads to estimated uncertainties that are too large to extract a numerically significant value for $\partial^{2}_{e}\langle z_{(1)}\rangle(p,e=0)$. We therefore construct a new grid near $e=0$ with uniform spacing in $e^{2}$ at fixed $p$ and recompute $\partial_{e^{2}}\langle z_{(1)}\rangle$ at $e=0$ using finite differencing.
Figure 1: Comparison of the analytical 10PN expression for $\langle z_{1}\rangle$, expanded in terms of $p$ and $e$, against our numerical results. On the horizontal axis we use the semi-latus rectum distance from the separatrix at $p_{\star}=6+2e$. The irregularities on the right side of the plot are due to numerical errors; in the large-$p$ region, we expect the PN expansion to be more accurate than the numerical results.
We have compared the 10PN redshift $\langle z_{(1)}\rangle^{\mathrm{10PN}}$ in terms of $(p,e)$, as well as its partial derivatives with respect to $p$ and $e$, with the corresponding numerical data, and plotted the relative difference in Fig. 1. Notably, we find the 10PN expansion agrees with the numerical results to at least 3 digits for all $p\gtrsim p_{\star}+2$. We observe only very mild loss of accuracy with increasing $e$ (though our comparison is limited to $e\leq 0.6$). At the top right of the plot, the disagreement is dominated by error in our numerical results rather than inaccuracy of the 10PN expansion. We have checked that analogous comparisons in terms of $(y,\lambda)$ lead to significantly worse agreement; we therefore omit the comparison.
IV.3 Constants of motion
We now extract the constants of motion from the redshift data, using Eq. (27). Since the derivatives of the redshift data are computed in terms of $(p,e)$, we use the chain rule to write
$\frac{\partial\langle z_{(1)}\rangle}{\partial\hat{\Omega}_{\phi}}$ $=\frac{\partial\langle z_{(1)}\rangle}{\partial p}\frac{\partial p}{\partial\hat{\Omega}_{\phi}}+\frac{\partial\langle z_{(1)}\rangle}{\partial e}\frac{\partial e}{\partial\hat{\Omega}_{\phi}}\,$, (54)
and similarly for $\Omega_{r}$. Since the matrix elements of $(\partial\hat{\Omega}_{A}/\partial\pi_{B})^{-1}$ are known exactly, we thus have everything we need to compute the constants of motion numerically.
However, we highlight that this approach suffers from large cancellations, causing significant loss of accuracy in our numerical results for the constants of motion relative to the underlying redshift data. To illustrate this, let us perform a numerical application of Eq. (54) for $p=115.76626305141966$ and $e=0.56454$ (i.e., $u=115/120$ and $w=98/100$); we find that
$0.29$ $=\underbrace{(-5.0243\times 10^{-4})\!\times\!(-9.2742\times 10^{6})}_{=456.96}$ (55)
$\qquad\qquad+\underbrace{(-9.6365\times 10^{-2})\!\times\!(4.8323\times 10^{5})}_{=-465.67}\,.$
Thus, the angular momentum loses $3$ digits of accuracy with respect to the numerical derivatives of the redshift. This feature is ubiquitous throughout the parameter space, particular for higher eccentricities $e\gtrsim 0.4$ and near the separatrix $\delta\lesssim 10^{-2}$ and at large separations $p\gtrsim 10^{2}$. Consequently, to compute first-order corrections $\hat{E}_{(1)}$, $\hat{L}_{(1)}$, and $\hat{J}_{r(1)}$ to $\sim 2$ digits of precision, we must accurately resolve derivatives with respect to $p$ and $e$ to $\sim 5$ digits.
Moreover, in the special case where $e=0$, we find that $\partial_{e}\langle z_{(1)}\rangle=0$ whereas $\partial e/\partial\hat{\Omega}_{\phi}$ and $\partial e/\partial\hat{\Omega}_{r}$ diverge as $1/e$ when $e\rightarrow 0$. Thus, to compute the 1SF corrections to the constants of motion for $e=0$ using this method, we need to control the second derivative $\partial_{e}^{2}\langle z_{(1)}\rangle(p,e=0)$, as discussed in the previous section. Further details of our numerical approach in the circular limit is provided in Appendices B and C. For this work, we accept that these cancellations and singular limits are a limitation of our $(p,e)$ parametrization. Future work will investigate alternative parametrizations that mitigate these issues.
The ancillary file data_generic.csv [1] provides, for each datapoint, the values of the following quantities: $p$, $e$, $\hat{\Omega}_{r}$, $\hat{\Omega}_{\phi}$, $y$, $\lambda$, $\langle z_{(0)}\rangle$, $\hat{E}_{(0)}$, $\hat{L}_{(0)}$, $\hat{J}_{r(0)}$, $\langle z_{(1)}\rangle$, $\hat{E}_{(1)}$, $\hat{L}_{(1)}$, and $\hat{J}_{r(1)}$. Absolute error estimates are also provided.
Following the same procedure, we also compute the constants of motion as PN expansions, starting from the 10PN expression for the 1SF redshift $\langle z_{(1)}\rangle$. This can be done analytically, and without any eccentricity re-expansion, thanks to the PN expansion of the Jacobian $(\partial\hat{\Omega}_{A}/\partial\pi_{B})$. Note that at leading PN order, one has
$\frac{\partial\langle z_{(1)}\rangle}{\partial p}\frac{\partial p}{\partial\hat{\Omega}_{r}}$ $=-\frac{\partial\langle z_{(1)}\rangle}{\partial e}\frac{\partial e}{\partial\hat{\Omega}_{r}}=-\frac{p^{3/2}}{3\sqrt{1-e^{2}}}+\mathcal{O}(\sqrt{p})\,$, (56a)
such that ${\partial\langle z_{(1)}\rangle}/{\partial\hat{\Omega}_{r}}=\mathcal{O}(\sqrt{p})$ instead of the $\mathcal{O}(p^{3/2})$ scaling one would naively expect. Even more remarkably,
$\frac{\partial\langle z_{(1)}\rangle}{\partial p}\frac{\partial p}{\partial\hat{\Omega}_{\phi}}$ $=-\frac{\partial\langle z_{(1)}\rangle}{\partial e}\frac{\partial e}{\partial\hat{\Omega}_{\phi}}$
$=\frac{p^{3/2}}{3\sqrt{1-e^{2}}}-\frac{(33+5e^{2})\sqrt{p}}{122\sqrt{1-e^{2}}}+\mathcal{O}\!\left(\frac{1}{\sqrt{p}}\right),$ (56b)
such that ${\partial\langle z_{(1)}\rangle}/{\partial\hat{\Omega}_{\phi}}=\mathcal{O}(1/\sqrt{p})$, which is two orders smaller than the $\mathcal{O}(p^{3/2})$ scaling one would naively expect. Consequently, our relative 10PN expression for the redshift can only give us access to the constants of motion at 9PN. These are given in terms of $(p,e)$ as
$\hat{E}_{(1)}$ $=\frac{1-e^{2}}{6p}\Bigg\{1-\frac{3(1-e^{2})}{p}+\mathcal{O}\!\left(\frac{1}{p^{2}}\right)\Bigg\}\,$, (57a)
$\hat{L}_{(1)}$ $=\frac{1}{\sqrt{p}}\Bigg\{-\frac{17}{12}+\frac{e^{2}}{6}+\mathcal{O}\!\left(\frac{1}{p}\right)\Bigg\}\,$, (57b)
$\hat{J}_{r(1)}$ $=-\frac{1}{3}\sqrt{\frac{p}{1-e^{2}}}\Bigg\{1+\mathcal{O}\!\left(\frac{1}{p}\right)\Bigg\}\,$, (57c)
and in terms of $(y,\lambda)$ as
$\hat{E}_{(1)}$ $=\frac{y}{6}\Bigg\{1+y\left(-1-\frac{2}{\lambda}\right)+\mathcal{O}(y^{2})\Bigg\},$ (58a)
$\hat{L}_{(1)}$ $=\sqrt{\frac{\lambda}{y}}\Bigg\{-\frac{1}{6}-\frac{5}{4\lambda}+\mathcal{O}(y)\Bigg\},$ (58b)
$\hat{J}_{r(1)}$ $=-\frac{1}{3\sqrt{y}}\Bigg\{1+\mathcal{O}(y)\Bigg\}.$ (58c)
The complete 9PN expressions, both in terms of $(p,e)$ and $(y,\lambda)$, are provided in the ancillary file PN_expressions.csv [1].
Figure 2: Comparison between our analytical (9PN) and numerical results for the 1SF corrections to the energy (top), angular momentum (middle), and radial action (bottom).
In Fig. 2, we plot the relative differences between our derived 9PN expressions and our numerical results for $\hat{E}_{(1)}$, $\hat{L}_{(1)}$, and $\hat{J}_{r(1)}$. We find relative errors $\lesssim 10^{-2}$ for $p\gtrsim p_{\star}+1$ in the case of $\hat{E}_{(1)}$, for $p\gtrsim p_{\star}+2$ in the case of $\hat{L}_{(1)}$, and for $p\gtrsim p_{\star}+3$ in the case of $\hat{J}_{r(1)}$. While the agreement generally improves at larger $p$, the relative errors start to grow again for $p\gtrsim 100$ and $e\gtrsim 0.2$. Like in the case of the redshift, this is due to numerical error in our data rather than a loss of accuracy in the PN expressions. The numerical error here is significantly larger than for the redshift due to the large cancellations described in Sec. IV.3. When taking these errors into account, the PN results for $\hat{E}_{(1)}$, $\hat{L}_{(1)}$, and $\hat{J}_{(r1)}$ lie within the estimated error bars of our numerical results for $p_{\star}\geq 2.7904, pp$$p_{\star}\geq 3.7861$, and $p-p_{\star}\geq 4.1915$, respectively. We have checked that analogous comparisons in terms of $(y,\lambda)$ yield significantly worse agreement.
Our use of a fixed-frequencies expansion also has another downside, besides numerical cancellations: $\hat{\Omega}_{A}$ are not valid phase-space coordinates along the isofrequency degeneracy curve discussed in Sec. III.1 (though they are valid coordinates in each of the two regions to the left and right of that curve). Because the inverse Jacobian $(\partial\hat{\Omega}_{A}/\partial\pi_{B})^{-1}$ diverges along that curve, the 1SF corrections to the constants of motion become ill-defined there in a fixed-frequencies expansion. This is illustrated in Fig. 3, where we see the correction $\hat{E}_{(1)}$ diverges in the approach to the isofrequency degeneracy.
Figure 3: The 1SF correction to the energy for $e=0.31974$ as a function of $p-p_{\star}$. We see that $\hat{E}_{(1)}$ diverges when it crosses the isofrequency curve at $p-p_{\star}=p_{\rm iso}-p_{\star}\approx 0.0023852$ (vertical dashed line). If we instead divide by the pole $(p-p_{\rm iso})^{-1}$, we get a smooth result.
V Applications
In this section we discuss the utility of the constants of motion in identifying special locations in the parameter space and in waveform generation. Specifically, we compute the location of circular orbits in the $(p,e)$ plane; the correction to the constants of motion along the separatrix; and corrections at the innermost stable circular orbit (ISCO). We then highlight the simplifications to waveform generation that arise from using balance laws for $E$ and $L$.
V.1 Circular orbits
Given the freedom in defining $(p,e)$ in the perturbed system, it is not guaranteed that $e=0$ will continue to specify a circular orbit. It is equally unclear what value $e$ takes in genuine quasicircular inspirals, which are characterized by the metric’s dependence on only one phase variable and frequency, $(\phi_{p},\Omega)$, rather than two [117].
We can define a (stable) circular orbit by a vanishing radial action, namely $\hat{J}_{r}=\hat{J}_{r(0)}+\varepsilon\,\hat{J}_{r(1)}=0$. This condition is preserved throughout a 1PA quasicircular evolution: from the evolution equation (320) in Ref. [107], one can easily show $\mathrm{d}J_{r}/\mathrm{d}t=0$ in the quasicircular case, due to the absence of any dependence on the radial phase $\varphi_{r}$. Expressed in terms of $(p,e)$, $\hat{J}_{r}=0$ determines a curve $e_{\odot}(p;\varepsilon)$, along which any circular (or quasicircular) orbit must lie. However, it will be more useful to work with the square of $e$, namely $e^{2}_{\odot}(p;\varepsilon)$, as we explain below.
Of course, since $e^{2}_{\odot}=0$ at geodesic order, $e^{2}_{\odot}(p;\varepsilon)$ must vanish when $\varepsilon\to 0$, but thanks to our expression for the radial action, we can now control the nonvanishing 1SF term in $e^{2}_{\odot}$. We obtain the correction numerically by injecting the ansatz
$e^{2}_{\odot}(p;\varepsilon)=\varepsilon\,\delta e_{\odot}^{2}(p)+\mathcal{O}(\varepsilon^{2})$ (59)
into the radial action. After replacement, we are left to solve
$\hat{J}_{r(0)}(p,0)+\varepsilon\,\delta e_{\odot}^{2}(p)\frac{\partial\hat{J}_{r(0)}}{\partial e^{2}}(p,0)+\varepsilon\,\hat{J}_{r(1)}(p,0)=\mathcal{O}(\varepsilon^{2})\,.$ (60)
The zeroth-order term vanishes by construction of the quasi-Keplerian variables. The derivative appearing in the subleading term is given analytically by
$\left.\frac{\partial\hat{J}_{r(0)}}{\partial e^{2}}\right|_{p}(p,0)$ $=\frac{p^{3/2}\sqrt{p-6}}{2(p-2)\sqrt{p-3}}\,$, (61)
allowing us to immediately solve Eq. (60) to find
$\delta e_{\odot}^{2}(p)$ $=-\frac{2(p-2)\sqrt{p-3}}{p^{3/2}\sqrt{p-6}}\hat{J}_{r(1)}(p,0)\,.$ (62)
Here we see the motivation for working with $e^{2}_{\odot}$ rather than $_{\odot}: ee_{\odot}$ scales with $\sqrt{\varepsilon}$ rather than $\varepsilon$. This is a consequence of the small-$e$ behavior of $\hat{J}_{r(0)}$, which has no linear term in a small-$e$ expansion.
Although Eq. (62) is exact, it suffers from hidden divergences in $\hat{J}_{r(1)}(p,0)$. To make these explicit, we build on previous works [16, 30] which have obtained the circular link between $x=[(m_{1}+m_{2})\Omega_{\phi}]^{2/3}$ and $W=(\Omega_{r}/\Omega_{\phi})^{2}$. The link reads $W^{\odot}(x)=1-6x+\varepsilon\rho(x)+\mathcal{O}(\varepsilon^{2})$, where $\rho(x)$ is provided numerically, is smooth in the domain $0<x\leq 1/6$, and converges to a finite value $\rho(1/6)\approx 0.83413$ [16]. In Appendix B, we show that $\hat{J}_{r(1)}(p,0)$ can be related to the $\rho(x)$ function via
$\hat{J}_{r(1)}(p,0)=\frac{p^{3/2}(p-6)^{3/2}\bigl[p\,\rho(1/p)-4\bigr]}{3\sqrt{p-3}(86-39p+4p^{2})}\,.$ (63)
Thus, we find that the eccentricity for circular orbits can be reexpressed as
$\delta e_{\odot}^{2}(p)=\ \frac{2(p-6)(p-2)}{3(86-39p+4p^{2})}\bigl[4-p\,\rho(1/p)\bigr]\,$, (64)
where the singular behavior is now fully explicit. We see that $\delta e_{\odot}^{2}(p)$ has a pole for $p_{\text{iso}}=(39+\sqrt{145})/8\approx 6.38$, which is exactly the value of $p$ corresponding to $e=0$ on the (geodesic) isofrequency curve [163]. Moreover, $\delta e_{\odot}^{2}(p)$ passes through zero at $p=p_{\text{sign}}\approx 6.95$, where $p_{\text{sign}}$ is the solution to $4-p\,\rho(1/p)=0$. This means that $\delta e_{\odot}^{2}$ changes sign at $p_{\text{sign}}$, becoming negative for $p_{\text{iso}}<p<p_{\text{sign}}$ and $0<\varepsilon\ll 1$, such that $e_{\odot}(p)$ become purely imaginary in that range. This points to the occasionally counterintuitive nature of $e$ in the fixed-frequencies gauge. For clarity, we have plotted the behavior of $\delta e_{\odot}^{2}(p)$ in Fig. 4.
Given the expression (36) for the orbital frequencies, we are now able to immediately find the frequencies for circular orbits. Injecting Eq. (64) into Eq. (36) and expanding to 1SF order, we find
$\hat{\Omega}_{\phi}^{\odot}(p)$ $=\frac{1}{p^{3/2}}+\varepsilon\,\frac{(22-10p+p^{2})}{p^{3/2}(86-39p+4p^{2})}\bigl[p\,\rho\left(1/p\right)-4\bigr],$ (65a)
$\hat{\Omega}_{r}^{\odot}(p)$ $=\frac{\sqrt{p-6}}{p^{2}}$
$\ +\varepsilon\,\frac{(-266+165p-32p^{2}+2p^{3})}{2p^{2}(86-39p+4p^{2})\sqrt{p-6}}\bigl[p\,\rho\left(1/p\right)-4\bigr].$ (65b)
Notice that the leading-order term in $\hat{\Omega}_{r}^{\odot}(p)$ vanishes for $p=6$, whereas the subleading term is divergent, which indicates a breakdown of the perturbative expansion around that point. We therefore will find it much more fruitful to work with the square of the radial frequency,
$(\hat{\Omega}_{r}^{\odot})^{2}(p)=\frac{{p-6}}{p^{4}}$
$\qquad\qquad+\varepsilon\,\frac{(-266+165p-32p^{2}+2p^{3})}{p^{4}(86-39p+4p^{2})}\bigl[p\,\rho\left(1/p\right)-4\bigr],$ (66)
where the subleading term is finite for $p=6$. Moreover, from Eq. (65a), one immediately deduces
$y_{\odot}(p)=\frac{1}{p}+\varepsilon\,\frac{2(22-10p+p^{2})}{3p(86-39p+4p^{2})}\bigl[p\,\rho\left(1/p\right)-4\bigr]\,.$ (67)
The expression for $\lambda_{\odot}(p)=\frac{3y_{\odot}}{\hat{\Omega}_{\phi}^{\odot}/\hat{\Omega}_{r}^{\odot}-1}$ is straighforward to obtain but too long to be worth displaying here.
In the spirit of the fixed frequency decomposition, we perform the following expansion
$z^{\odot}(y)$ $=z_{(0)}^{\odot}(y)+\varepsilon\,z_{(1)}^{\odot}(y)+\mathcal{O}(\varepsilon^{2})\,$, (68a)
$\hat{E}^{\odot}(y)$ $=\hat{E}_{(0)}^{\odot}(y)+\varepsilon\,\hat{E}_{(1)}^{\odot}(y)+\mathcal{O}(\varepsilon^{2})\,$, (68b)
$\hat{L}^{\odot}(y)$ $=\hat{L}_{(0)}^{\odot}(y)+\varepsilon\,\hat{L}_{(1)}^{\odot}(y)+\mathcal{O}(\varepsilon^{2})\,$, (68c)
$\hat{\Omega}_{r}^{\odot}(y)$ $=\hat{\Omega}_{r(0)}^{\odot}(y)+\varepsilon\,\hat{\Omega}_{r(1)}^{\odot}(y)+\mathcal{O}(\varepsilon^{2})\,.$ (68d)
Note that, by definition, the radial action is vanishing to all orders for circular orbits. At geodesic order, we have the well known expressions
$z_{(0)}^{\odot}(y)$ $=\sqrt{1-3y}\,$, (69a)
$\hat{E}_{(0)}^{\odot}(y)$ $=\frac{1-2y}{\sqrt{1-3y}}\,$, (69b)
$\hat{L}_{(0)}^{\odot}(y)$ $=\frac{1}{\sqrt{y(1-3y)}}\,$, (69c)
$\hat{\Omega}_{r(0)}^{\odot}(y)$ $=y^{3/2}\sqrt{1-6y}\,.$ (69d)
At 1SF order, we need to invert the relation (65a) to read off
$\delta p_{\odot}(y)$ $=\frac{2(1-10y+22y^{2})}{3y^{2}(4-39y+86y^{2})}\big[\rho(y)-4y\big]\,$, (70a)
$\delta e^{2}_{\odot}(y)$ $=\frac{2(1-6y)(1-2y)}{3y(4-39y+86y^{2})}\big[4y-\rho(y)\big]\,.$ (70b)
The results are then obtained numerically and provided in the ancillary file data_circular.csv, which contains, for each datapoint, the values of the following quantities: $y$, $\hat{\Omega}_{\phi}$, $\hat{\Omega}_{r(0)}$, $\langle z_{(0)}^{\odot}\rangle$, $\hat{E}_{(0)}^{\odot}$, $\hat{L}_{(0)}^{\odot}$, $\hat{\Omega}_{r(1)}^{\odot}$, $\langle z_{(1)}^{\odot}\rangle$, $\hat{E}_{(1)}^{\odot}$, $\hat{L}_{(1)}^{\odot}$, $\delta p_{\odot}$, and $\delta e_{\odot}^{2}$. Absolute error estimates are also provided.
Figure 4: Plot of $\delta e_{\odot}^{2}(p)$ using various schemes. The blue dots and error bars are obtained by evaluating Eq. (62) using our numerical results for $\hat{J}_{r(1)}^{\odot}(p)$. The green dot-dashed line is obtained by replacing $\rho(x)$ by its 9PN expression in Eq. (64), whereas the orange dashed line is obtained by using a non-resummed 9PN expansion (71) for $\delta e_{\odot}^{2}(p)$. The vertical grey dotted lines correspond to, from left to right, $p=p_{\text{iso}}\approx 6.38$ and $p=p_{\text{sign}}\approx 6.95$. Thus, for $p_{\text{iso}}<p<p_{\text{sign}}$, we have $\delta e_{\odot}^{2}(p)<0$ and $e_{\odot}$ is an imaginary number.
We now proceed to determine these circular links as PN expansions. First, we compute the self-force corrections to the circular squared eccentricity $e_{\odot}^{2}$ in terms of $p$. Using our 9PN expansion of $\hat{J}_{r(1)}(p,0)$, or equivalently, using the PN expansion for $\rho(x)$ provided in Eq. (9) of Ref. [30], we can approximate the location of circular orbits as
$e^{2}_{\odot}$ $=\varepsilon\,\Bigg\{\frac{2}{3}-\frac{7}{6p}+\mathcal{O}\left(\frac{1}{p^{2}}\right)\Bigg\}+\mathcal{O}(\varepsilon^{2})\,$, (71)
where the complete 9PN expression is given in the ancillary file PN_expressions.csv [1].
Similarly, it is straightforward to PN-expand the frequencies $\hat{\Omega}_{\phi}^{\odot}(p)$ and $\hat{\Omega}_{r}^{\odot}(p)$ provided in Eqs. (65); then, eliminating $p$, we can find the gauge-invariant relation between the two frequencies. The latter is most conveniently represented through the fixed-frequency expansion of $\lambda^{\odot}(y)=\lambda_{(0)}^{\odot}+\varepsilon\,\lambda_{(1)}^{\odot}+\mathcal{O}(\varepsilon^{2})$; we find that
$\lambda_{(0)}^{\odot}(y)$ $=1-\frac{9}{2}y-\frac{9}{4}y^{2}+\mathcal{O}(y^{3})\,$, (72a)
$\lambda_{(1)}^{\odot}(y)$ $=-\frac{2}{3}+\frac{7}{3}y+y^{2}\left(\frac{379}{12}-\frac{41}{32}\pi^{2}\!\right)+\mathcal{O}(y^{3})\,$, (72b)
where the complete 9PN expressions are given in the ancillary file PN_expressions.csv [1]. After appropriate rescaling, we find agreement at geodesic and 1SF orders with the 3PN-accurate expression (7.4a) of Ref. [152], which includes contributions from the tail that enters the 4PN dynamics.
It is possible to deduce other gauge-invariant links for circular orbits. For instance, the energy is related to the angular momentum through $\hat{E}^{\odot}(\hat{L})=\hat{E}_{(0)}^{\odot}(\hat{L})+\varepsilon\,\hat{E}_{(1)}^{\odot}(\hat{L})+\mathcal{O}(\varepsilon^{2})$. We find that
[…]
where the complete 9PN expressions are given in the ancillary file PN_expressions.csv [1]. After rescaling by the correct prefactor, we found perfect agreement with the 4PN circular link in Eq. (5.3) of Ref. [56] at geodesic and 1SF orders.
V.2 Homoclinic orbits on the separatrix
We now turn our attention to the separatrix between bound and plunging orbits. Orbits on this separatrix correspond to homoclinic trajectories that are maximally “zoom-whirly”, with a single radial excursion (the “zoom”) to apastron and an infinite “whirl” phase in approaching and departing periastron [40, 63, 106, 126]; a point with orbital radius precisely at this periastron is a fixed point, corresponding to an unstable circular orbit. The homoclinic orbits necessarily have an infinite radial period, taking infinite time to leave the fixed point in the past and infinite time to reach it in the future. Hence, they have vanishing radial frequency, and we can use $\Omega_{r}=0$ to define the location of the separatrix, to all orders in the mass ratio.
At geodesic order, the condition $\Omega_{r}=0$ corresponds to $p=6+2e$. Since we have chosen the quasi-Keplerian parameters $(p,e)$ to be geodesically related to the frequencies, the separatrix remains at $p=6+2e$ at all SF orders. Denoting points on the separatrix with a star, as in $(p_{\star},e_{\star})$, we parameterize the separatrix location as $p_{\star}(e)=6+2e$. The azimuthal frequency can then be obtained as a pure function of the eccentricity, $\Omega_{\phi}^{\star}(e)=\Omega_{\phi}(p_{\star}(e),e)$. Computing this explicitly from Eq. (36) is in fact nontrivial because the limit of $\Omega_{\phi}(p,e)$ as $\delta\equiv p-p_{\star}(e)\rightarrow 0$ is singular. To examine this limit, we first apply the relations between elliptic functions provided in Eqs. (114)–(115), then use the near-separatrix behavior of the elliptic integral function of the third kind $\Pi(\cdot|\cdot)$ provided in Eq. (3.5) of Ref. [108], as well as the limits
$K\left(\frac{1}{1+z}\right)$ $\underset{z\rightarrow 0^{+}}{\sim}-\frac{1}{2}\ln\left(\frac{z}{16}\right)\,$, (74a)
$E\left(\frac{1}{1+z}\right)$ $\underset{z\rightarrow 0^{+}}{\longrightarrow}1\,.$ (74b)
Using the expansions of the elliptic integrals, we can determine that all three periods diverge logarithmically:
$T_{r}$ $\underset{\delta\rightarrow 0^{+}}{\sim}\frac{4(3+e)^{2}}{\sqrt{e}(1+e)^{3/2}}\ln\left(\frac{1}{\delta}\right)\,$, (75a)
$\Phi$ $\underset{\delta\rightarrow 0^{+}}{\sim}\sqrt{\frac{2(3+e)}{e}}\ln\left(\frac{1}{\delta}\right)\,$, (75b)
$\mathcal{T}_{r}$ $\underset{\delta\rightarrow 0^{+}}{\sim}\sqrt{\frac{8(3-e)(3+e)^{3}}{e(1+e)^{3}}}\ln\left(\frac{1}{\delta}\right)\,$, (75c)
implying that $\Omega_{r}$ vanishes in this limit (as expected). On the other hand, the geodesic redshift $\langle z_{0}\rangle=\hat{T}_{r}/\hat{\mathcal{T}}_{r}$ has a finite near-separatrix limit:
$\langle z_{(0)}^{\star}\rangle=\sqrt{\frac{3-e}{2(3+e)}}=\sqrt{\frac{6}{p}-\frac{1}{2}}=\sqrt{1-3y}\,.$ (76)
Finally, the azimuthal frequency $\hat{\Omega}_{\phi}=\Phi/\hat{\mathcal{T}}_{r}$ also has a finite limit on the separatrix [108]:
$y^{\star}=\left(\hat{\Omega}_{\phi}^{\star}\right)^{2/3}=\frac{1+e^{\star}}{6+2e_{\star}}=\frac{1}{2}-\frac{2}{p^{\star}}\,.$ (77)
This expression is identical to that of an unstable circular (geodesic) orbit: $y=1/\hat{r}+\mathcal{O}(\varepsilon)$, where the radius of the periastron $\hat{r}_{\rm min}=p/(1+e)$ is evaluated for $p=p_{\star}=6+2e$ on the separatrix. Finally, note that the vanishing of the radial frequency immediately leads to $\lambda^{\star}(e)=0$.
The geodesic energy, angular momentum, and radial action all admit a finite limit on the separatrix. The geodesic energy and angular momentum reduce to the usual geodesic expressions for circular orbits [48] with (effective) radius $\hat{r}_{\text{min}}$. In the case of the radial action, using the previously established limits for the elliptic integrals, we find a complicated expression that depends of Legendre’s elliptic integrals of the third kind $\Pi(\cdot|\cdot)$ as well as the functions $\Lambda(e)$ and $\Theta(e)$ introduced in Ref. [108]. Alternatively, by directly integrating Eq. (32) with $p=p_{\star}=6+2e$, we were able to find a more tractable expression [^8] This integral was performed in App. A of Ref. [126] for Kerr spacetime; one recovers the Schwarzschild result of Eq. (78c) by setting $a=0$., which agrees numerically with the former. The energy, angular momentum, and radial action on the separatrix finally read
$\hat{E}_{(0)}^{\star}$ $=\sqrt{\frac{8}{(3+e)(3-e)}}\,$, (78a)
$\hat{L}_{(0)}^{\star}$ ${}={}\frac{2(3+e)}{\sqrt{(1+e)(3-e)}}\,$, (78b)
$\hat{J}_{r{(0)}}^{\star}$ $=\frac{2}{\pi\sqrt{9-e^{2}}}\Bigg\{\frac{7+e^{2}}{\sqrt{1-e^{2}}}\arctan\left(\sqrt{\frac{2e}{1-e}}\right)$ (78c)
$\quad\quad-4\sqrt{2}\,\text{arctanh}\left(\sqrt{\frac{e}{1+e}}\right)-(3+e)\sqrt{\frac{2e}{1+e}}\Bigg\}\,.$
It is immediate to re-express these in terms of either $p$ or $y$ using Eq. (77).
We now use the results of this paper to compute the self-force corrections to the redshift, energy, angular momentum, and radial action on the separatrix. These invariants can be decomposed as usual into a (fixed-frequencies) expansion in the mass ratio:
$\langle z^{\star}\rangle(e)$ $=\langle z_{(0)}^{\star}\rangle(e)+\varepsilon\,\langle z_{(1)}^{\star}\rangle(e)+\mathcal{O}(\varepsilon^{2})\,$, (79a)
$\hat{E}^{\star}(e)$ $=\hat{E}_{(0)}^{\star}(e)+\varepsilon\,\hat{E}_{(1)}^{\star}(e)+\mathcal{O}(\varepsilon^{2})\,$, (79b)
$\hat{L}^{\star}(e)$ $=\hat{L}_{(0)}^{\star}(e)+\varepsilon\,\hat{L}_{(1)}^{\star}(e)+\mathcal{O}(\varepsilon^{2})\,$, (79c)
$\hat{J}_{r}^{\star}(e)$ $=\hat{J}_{r(0)}^{\star}(e)+\varepsilon\,\hat{J}_{r(1)}^{\star}(e)+\mathcal{O}(\varepsilon^{2})\,.$ (79d)
The 1SF corrections are estimated by fitting for the first-order terms in (79) using our values for $\langle z_{(1)}\rangle$, $\hat{E}_{(1)}$, $\hat{L}_{(1)}$, and $\hat{J}_{r(1)}$ near the separatrix. These values are reported in the ancillary file data_separatrix.csv [1], which includes the following quantities: $p$, $e$, $\hat{\Omega}_{\phi}$, $y$, $\langle z_{(0)}^{\star}\rangle$, $\hat{E}_{(0)}^{\star}$, $\hat{L}_{(0)}^{\star}$, $\hat{J}_{r(0)}^{\star}$, $\langle z_{(1)}^{\star}\rangle$, $\hat{E}_{(1)}^{\star}$, $\hat{L}_{(1)}^{\star}$, and $\hat{J}_{r(1)}^{\star}$. Absolute error estimates are also provided.
Finally, we note that in an effective-potential description, points on the separatrix correspond to orbits whose energies are at the peak of the effective potential. That textbook description of the geodesic dynamics extends straightforwardly to 1SF order: The condition $g^{\alpha\beta}p_{\alpha}p_{\beta}=(m_{2})^{2}$, together with $p_{t}=E+\varepsilon\delta p_{t}$ and $p_{\phi}=L+\varepsilon\delta p_{\phi}$, implies the (adimensionalized) radial equation
$\left(\frac{d\hat{r}_{p}}{d\hat{\tau}}\right)^{\!2}=\hat{E}^{2}-V(\hat{r}_{p},\hat{E},\hat{L},\varepsilon),$ (80)
with the effective potential
$V=f_{p}\left(1+\frac{\hat{L}^{2}}{\hat{r}_{p}^{2}}\right)+2\varepsilon\left(f_{p}\frac{\hat{L}\delta\hat{p}_{\phi}}{\hat{r}_{p}^{2}}-\hat{E}\delta\hat{p}_{t}\right).$ (81)
[…]
Just as in the usual geodesic description, Eq. (80) implies
$\hat{E}^{2}=V(\hat{r}_{\rm min/max},\hat{E},\hat{L},\varepsilon)$ (82)
at turning points and on circular orbits. Differentiating Eq. (80) with respect to $\hat{\tau}$ and restricting to the conservative sector (where $E$ and $L$ are constants), we also get
$\frac{d^{2}\hat{r}_{p}}{d\hat{\tau}^{2}}=-\frac{1}{2}\frac{\partial V}{\partial\hat{r}_{p}},$ (83)
again just as in the geodesic case. Unstable circular orbits correspond to maxima of the effective potential, where $\frac{\partial V}{\partial\hat{r}_{p}}=0$ picks out circular orbits and $\frac{\partial^{2}V}{\partial r_{p}^{2}}<0$ ensures they are unstable. The two relations (82) and $\frac{\partial V}{\partial\hat{r}_{p}}=0$, with $\frac{\partial^{2}V}{\partial r_{p}^{2}}<0$, determine $\hat{E}_{\rm UCO}(\hat{L},\varepsilon)$ and $\hat{r}_{\rm UCO}(\hat{L},\varepsilon)$. The homoclinic orbits are eccentric trajectories with $\hat{E}^{\star}(\hat{L},\varepsilon)=\hat{E}_{\rm UCO}(\hat{L},\varepsilon)$ and $\hat{r}_{\rm min}(\hat{L},\varepsilon)=\hat{r}_{\rm UCO}(\hat{L},\varepsilon)$, such that they precisely reach the peak of the effective potential at their periapsis, in the infinite past and future.
[…]
V.3 Innermost stable circular orbit
We now specialize to the circular case and verify that we are able to recover known results for the shift in the frequency of the innermost stable circular orbit (ISCO), as well as the values of constants of motion at the ISCO. Most of these quantities are highly singular near the ISCO: limits do not necessarily commute, Taylor-expansions are often singular when expanding around $(p,e)=(6,0)$ and different orders of operations can yield different results. Thus, we here follow the usual definition in the literature: we first restrict to the curve of circular orbits, then take the limit towards the ISCO. Precisely at the ISCO, the curve of circular orbits (defined by $J_{r}=0$) intersects the separatrix (defined by $\Omega_{r}=0$), and we can equivalently define the ISCO by $\Omega_{r}=0=J_{r}$.
At geodesic order, $(p,e)=(6,0)$ and Eq. (77) immediately yields $y_{\text{ISCO}}=1/6+\mathcal{O}(\varepsilon)$. The redshift, energy, angular momentum, and radial action read, respectively, $z_{(0)}^{\text{ISCO}}=1/\sqrt{2}$, $\hat{E}_{(0)}^{\text{ISCO}}=2\sqrt{2}/3$, $\hat{L}_{(0)}^{\text{ISCO}}=2\sqrt{2}$ and $\hat{J}_{r(0)}^{\text{ISCO}}=0$.
To extend these results to 1SF order, we first want to find the values of $(p,e)$ at the ISCO, which we denote $(p^{\text{ISCO}},e^{\text{ISCO}})$. We will first restrict to the curve of circular orbits ($\hat{J}_{r}=0$), such that $p$ and $e$ are related by Eqs. (59)–(64). On that curve, the radial frequency is a function only of $p$, denoted $\Omega_{r}^{\odot}(p)$; see (65). To find the ISCO, which is the marginally stable circular orbit, we must find $p$ such that $\Omega_{r}^{\odot}=0$. However, we must account for the fact that our expansion for $\Omega_{r}^{\odot}(p)$ is ill-behaved around $p=6$: the geodesic term in Eq. (65) scales as $\sqrt{p-6}$, such that only the right-handed limit is physically well defined and vanishing, whereas the subleading term scales as $1/\sqrt{p-6}$ and thus blows up. Since we are working perturbatively and performing a Taylor expansion, we want to work with regular quantities. Fortunately, we have found in Eq. (V.1) that $(\Omega_{r}^{\odot})^{2}$ is regular around $p=6$, so we will instead solve for $(\Omega_{r}^{\odot})^{2}=0$. This choice is in line with Refs. [16, 18, 19], which solve for $W=\Omega_{r}^{2}/\Omega_{\phi}^{2}=0$. To do this in practice, we write the following ansatz for the value of the semilatus rectum on the ISCO:
$p_{\text{ISCO}}=6+\,\varepsilon\,\delta p_{\text{ISCO}}+\mathcal{O}(\varepsilon^{2}).$ (85)
Plugging the ansatz for $p_{\text{ISCO}}$ into Eq. (V.1), reexpanding in $\varepsilon$ and solving for $p_{\text{ISCO}}$, we find that
$p_{\mathrm{ISCO}}$ $=6+2\varepsilon\bigl[3\,\rho(1/6)-2\bigr]+\mathcal{O}(\varepsilon^{2})\,.$ (86)
Injecting this value into Eq. (67), we finally find that the frequency at the ISCO reads
$y_{\text{ISCO}}$ $=\frac{1}{6}+\varepsilon\,\frac{3\,\rho(1/6)-2}{18}+\mathcal{O}(\varepsilon^{2})\,.$ (87)
This is of course in agreement with the expression we would have found by directly solving $W=(\Omega_{r}^{\odot}/\Omega_{\phi}^{\odot})^{2}=0$ in Eq. (B).
Finally, we compute the 1SF corrections to the constants of motion at the ISCO. We inject the expression (87) for $y_{\text{ISCO}}$ into Eqs. (68)–(69) and expand in small $\varepsilon$. We find that
$z^{\text{ISCO}}$ $=\frac{1}{\sqrt{2}}+\varepsilon\,z_{(1)}^{\text{ISCO}}+\mathcal{O}(\varepsilon^{2})\,$, (88a)
$\hat{E}^{\text{ISCO}}$ $=\frac{2\sqrt{2}}{3}+\varepsilon\,\hat{E}_{(1)}^{\text{ISCO}}+\mathcal{O}(\varepsilon^{2})\,$, (88b)
$\hat{L}^{\text{ISCO}}$ $=2\sqrt{3}+\varepsilon\,\hat{L}_{(1)}^{\text{ISCO}}+\mathcal{O}(\varepsilon^{2})\,$, (88c)
where $\hat{E}_{(1)}^{\text{ISCO}}=\hat{E}_{(1)}^{\odot}(6)$, $\hat{L}_{(1)}^{\text{ISCO}}=\hat{L}^{\odot}_{(1)}(6)$, and
$z_{(1)}^{\text{ISCO}}=\frac{1}{6\sqrt{2}}\left[2-3\rho\left(\frac{1}{6}\right)\right]+z_{(1)}^{\odot}(6)\,.$ (89)
By definition, the radial action is vanishing to all orders on the ISCO.
To numerically evaluate the ISCO frequency shift, we obtain $\rho(1/6)$ by computing $\partial_{p}\langle z_{(1)}\rangle(p,0)$, $\partial_{e}^{2}\langle z_{(1)}\rangle(p,0)$, and $\rho(1/p)$ near the separatrix [using Eq. (63)] and fit for its value on the ISCO; see Appendix C for further details. From this procedure we obtain
$\rho(1/6)=0.8340(2)\,.$ (90)
Decomposing the reduced ISCO frequency in powers of the mass ratio, namely as
$\hat{\Omega}_{\text{ISCO}}=\hat{\Omega}_{(0)}^{\text{ISCO}}+\varepsilon\,\hat{\Omega}_{(1)}^{\text{ISCO}}+\mathcal{O}(\varepsilon^{2}),$ (91)
we then find that the relative ISCO frequency shift reads
$\frac{\hat{\Omega}_{(1)}^{\text{ISCO}}}{\hat{\Omega}_{(0)}^{\text{ISCO}}}=\frac{3}{2}\,\rho(1/6)-1=0.2510(2)\,$, (92)
which is in agreement [^9] The value reported in Ref. [93] is $C_{\Omega}-1=\hat{\Omega}^{\mathrm{ISCO}}_{(1)}/\hat{\Omega}^{\mathrm{ISCO}}_{(0)}=0.25101539(4)$. with the high-accuracy calculations reported in Ref. [93]. Similarly to $\rho$, we follow the methods described in Appendix C to fit for the near-ISCO behavior, leading to
$z_{(1)}^{\mathrm{ISCO}}$ $=+0.08885(6)\,$, (93a)
$\hat{E}_{(1)}^{\mathrm{ISCO}}$ $=+0.019424700(6)\,$, (93b)
$\hat{L}_{(1)}^{\mathrm{ISCO}}$ $=-0.80219089(3)\,.$ (93c)
After properly accounting for the different conventions, we find very good agreement with the highly accurate redshift for circular orbits used in the 1PA waveform model [167]; good agreement is also found with Refs. [99, 104]
V.4 Waveform generation
The most significant application of the conserved quantities is in waveform models that employ balance laws, which express the binary evolution in terms of fluxes of energy and angular momentum out to infinity and into the primary black hole. Here we outline the construction of such models at 1PA order. We also explain the critical assumption they would currently need to rely on.
As described in Sec. II, the general framework for a 1PA waveform model is well understood. If the waveform is decomposed in spin-weighted harmonics, as in
$h_{+}-ih_{\times}=\frac{m_{1}}{r}\sum_{\ell m}h_{\ell m}(u){}_{-2}Y_{\ell m}(\theta,\phi),$ (94)
then the expansions (9) and (10) imply that each $(\ell,m)$ mode of the waveform is given by
$h_{\ell m}=\sum_{n=-\infty}^{\infty}\Bigl[\varepsilon h^{(1)}_{\ell mn}(\pi_{A})+\varepsilon^{2}h^{(2)}_{\ell mn}(\pi_{A})\\ +{\cal O}(\varepsilon^{3})\Bigr]e^{-i(m\varphi_{\phi}+n\varphi_{r})},$ (95)
where we have adopted the mode labels $k_{A}=(n,m)$ common in the literature [49]. The time evolution of the phase-space coordinates $(\varphi_{A},\pi_{A})$, and hence the waveform’s time dependence, is governed by the ordinary differential equations (4) and (5). In those evolution equations, the 0PA dynamics are well understood and efficiently evaluated in terms of gravitational-wave fluxes [142, 90]. But historically, the 1PA terms $\Omega^{(1)}_{A}$ and $F_{A}^{(1)}$ have only been written in terms of the local self-forces $f^{\alpha}_{(1)}$ and $f^{\alpha}_{(2)}$ acting on the particle [130, 113] (or equivalent local expressions in terms of a pseudo-Hamiltonian [107]).
In Refs. [107, 121] (building on Ref. [76]), we emphasized that the conservative sector of the complete 1PA dynamics, as represented by the frequency correction $\Omega^{(1)}_{A}$, can be most straightforwardly computed in terms of the redshift $\langle z_{(1)}\rangle$. However, this leaves the dissipative 1PA term, $F^{(1)}_{A}$, as a complicated mix of local 1SF and 2SF quantities; as an example, see Eq. (A10) of Ref. [117] for the simple case of quasicircular orbits. Even in that simple case, the local 1PA expressions have never been evaluated. Instead, the complicated local quantities are replaced by simple fluxes. Here we explain that replacement, generalizing the argument from the quasicircular case in Refs. [167, 9].
The argument begins from exact, fully nonlinear balance laws at infinity and at the primary’s horizon. At future null infinity, the Bondi mass and angular momentum, $P^{\infty}_{A}=(E^{\infty},L^{\infty})$, satisfy well-known, exact balance laws [53]:
$\frac{dP^{\infty}_{A}}{du}=-{\cal F}^{\infty}_{A}(u,\varepsilon),$ (96)
where ${\cal F}^{\infty}_{A}$ are the fluxes of energy and angular momentum to infinity. Similarly, at the primary black hole’s horizon, the black hole mass and spin, $P^{\rm BH}_{A}=(m_{1},s_{1})$, satisfy exact balance laws [15, 47]:
$\frac{dP^{\rm BH}_{A}}{dv}={\cal F}^{\cal H}_{A}(v,\varepsilon),$ (97)
where ${\cal F}^{\cal H}_{A}$ are the fluxes of energy and angular momentum down the horizon.
These exact balance laws for $P^{\infty}_{A}$ and $P^{\rm BH}_{A}$ translate into approximate ones for the orbital quantities $P_{A}$, subject to the following unproved assumption: on average, the total (Bondi) mass and angular momentum are equal, up to numerically small differences, to the binary’s total mechanical energy and angular momentum. Here by mechanical energy we mean the sum of the black hole’s mass and the orbital energy, and likewise for the mechanical angular momentum. In short, the assumption is
$P_{A}+\bigl\langle P^{\rm BH}_{A}\bigr\rangle=\bigl\langle P_{A}^{\infty}\bigr\rangle.$ (98)
This relationship is required to hold on each time slice $s=\text{constant}$, where we recall that $s$ reduces to $v$ at the horizon, $t$ at the particle, and $u$ at future null infinity. Taking a time derivative of Eq. (98) and appealing to the exact balance laws above, we obtain
$\frac{dP_{A}}{ds}=-{\cal F}_{A},$ (99)
where we have defined the total torus-averaged fluxes
${\cal F}_{A}:=\bigl\langle{\cal F}^{\infty}_{A}\bigr\rangle+\bigl\langle{\cal F}^{\cal H}_{A}\bigr\rangle.$ (100)
The relationship (98) holds exactly at 0PA order [69, 115, 160, 151]. However, at least for quasicircular orbits, it is known to fail at 1PA order [129], specifically at 4PN. In Ref. [154], one of us showed concretely that, after orbit-averaging, the Bondi mass differs from the mechanical energy by a 4PN Schott term related to dissipative tail effects. This will be extended to a fully relativistic result in Ref. [80], implying a corrected relationship
$P_{A}+\bigl\langle P^{\rm BH}_{A}\bigr\rangle+\delta P^{\rm Schott}_{A}=\bigl\langle P_{A}^{\infty}\bigr\rangle.$ (101)
In principle, an exact 1PA formula for $dP_{A}/ds$ can be obtained by differentiating this equality, but $\delta P^{\rm Schott}_{A}$, which depends on our choice of spacetime foliation, is not yet known for eccentric orbits. On the other hand, the results in Ref. [129] suggest that the Schott term is numerically small (in addition to carrying a relative factor of $\varepsilon$) and it has been neglected in all 1PA waveform models; see Ref. [9] for a tentative assessment of the impact of this small error on gravitational-wave phasing.
If we proceed cavalierly under the assumption that $\delta P^{\rm Schott}_{A}$ is numerically small, then the evolution equations (4) and (5) can be replaced with
$\frac{d\varphi_{A}}{dt}$ $=\Omega_{A}^{(0)}(\hat{P}_{B})+\varepsilon\Omega_{A}^{(1)}(\hat{P}_{B}),$ (102)
$\frac{d\hat{P}_{A}}{dt}$ $=-\varepsilon\left[{\cal F}^{(0)}_{A}(\hat{P}_{B})+\varepsilon{\cal F}^{(1)}_{A}(\hat{P}_{B})\right].$ (103)
Here, given the definition of $s$, we freely use $t$ at the particle with the understanding that it is replaced by $u$ in the waveform. The 0PA forcing function ${\cal F}^{(0)}_{A}(\hat{P}_{B})$ is the standard 0PA flux (normalized by $m_{2}$) computed from the first-order waveform mode amplitudes $h^{(1)}_{\ell mn}$ (at the horizon and infinity) [130], and ${\cal F}^{(1)}_{A}$ is computed from products of the amplitudes $h^{(1)}_{\ell mn}$ and $h^{(2)}_{\ell mn}$. [^10] In principle this includes 1PA fluxes at the primary’s horizon, but they have not yet been included in 1PA waveform models. We again refer to Ref. [9] for a discussion of how this omission impacts the waveform phase. The frequency correction $\Omega^{(1)}_{A}(\hat{P}_{B})$ is given by Eq. (46).
Even if the field equations are solved on a grid of $(p,e)$ values, one can still work directly with the evolution equations (102) and (103) by adopting the “fixed constants” gauge discussed in Sec. IV. In this gauge the quasi-Keplerian parameters, denoted $\pi^{\rm FC}_{A}$, are related to $(E,L)$ by the Schwarzschild-geodesic relationship: $P_{A}(\pi^{\rm FC}_{B},\varepsilon)=P^{(0)}_{A}(\pi^{\rm FC}_{B})$. Hence, while evolving $P_{A}$, one can immediately find stored data at corresponding values of $\pi^{\rm FC}_{A}$.
Alternatively, from Eq. (103) we can immediately derive evolution equations for our fixed-frequency variables; recall these satisfy $\Omega_{A}(\pi_{B},\varepsilon)=\Omega^{(0)}_{A}$. Given the chain rule $\dot{\pi}_{A}=(\partial\pi_{A}/\partial P_{B})\dot{P}_{B}$, we have
$\frac{d\varphi_{A}}{dt}$ $=\Omega_{A}^{(0)}(\pi_{B}),$ (104a)
$\frac{d\pi_{A}}{dt}$ $=-\varepsilon\frac{\partial\pi_{A}}{\partial P_{B}}\Bigl\{{\cal F}^{(0)}_{B}(\pi_{C})$
[…]
where the Jacobian is given by
$\frac{\partial\pi_{A}}{\partial P_{B}}$ $=\begin{pmatrix}\frac{\partial E}{\partial p}&\frac{\partial E}{\partial e}\\ \frac{\partial L}{\partial p}&\frac{\partial L}{\partial e}\end{pmatrix}^{\!-1}=\frac{1}{\mathrm{det}\left|\frac{\partial P_{A}}{\partial\pi_{B}}\right|}\!\begin{pmatrix}\frac{\partial L}{\partial p}&-\frac{\partial E}{\partial e}\\ -\frac{\partial L}{\partial p}&\frac{\partial E}{\partial e}\end{pmatrix}\!$. (105)
This matrix can be expanded for small $\varepsilon$ as
$\frac{\partial\pi_{A}}{\partial P_{B}}=\frac{\partial\pi_{A}}{\partial P^{(0)}_{B}}-\varepsilon\frac{\partial\pi_{A}}{\partial P^{(0)}_{C}}\frac{\partial P^{(1)}_{C}}{\partial\pi_{D}}\frac{\partial\pi_{D}}{\partial P^{(0)}_{B}}+{\cal O}(\varepsilon^{2}),$ (106)
where it is understood that $\partial\pi_{A}/\partial P^{(0)}_{B}$ represents the inverse of $\partial P^{(0)}_{B}/\partial\pi_{A}$. However, leaving the Jacobian in unexpanded form is likely to lead to a more accurate waveform model, in analogy with the “resummed” models for quasicircular binaries in Refs. [114, 88, 87]. More specifically, it is likely important to leave the determinant denominator in Eq. (105) intact, as it preserves the Jacobian’s singularity at the corrected separatrix rather than at the background one.
Regardless of whether the Jacobian $\partial\pi_{A}/\partial P_{B}$ is expanded, we note that since the corrections $P^{(1)}_{A}$ are already computed as first derivatives of the redshift, the Jacobian requires second derivatives of the redshift. In App. E, we derive a formula which can be used to express $\partial P_{A}^{(1)}/\partial\pi_{B}$ explicitly in terms of $\partial^{2}\langle z_{1}\rangle/(\partial\pi_{A}\partial\pi_{B})$.
VI Discussion and future applications
In this paper we have carried out the first comprehensive calculations of the energy, angular momentum, and radial action of nonspinning binaries at 1SF order beyond the geodesic approximation, providing analytical results up to 9PN and numerical results across the parameter space up to eccentricity $e=0.5$. These data and formulas are provided in the Supplemental Material [1]. Although, for simplicity, we have restricted our calculations to a nonspinning primary, the formalism [107] as well as the numerical [157, 121] and analytical [29, 119] methods are already well developed for eccentric orbits around a Kerr primary.
Our primary motivation for studying these constants of motion is described in Sec. V.4: to utilize them in 1PA waveform generation, capitalizing on the streamlined form that the waveform takes when written in terms of balance laws. Without access to the constants of motion, 1PA waveform models require calculation of local 2SF forces, which have never been carried out. Balance laws evade that requirement, as demonstrated in the quasicircular case [167].
Actually implementing the 1PA waveform-generation scheme outlined in Sec. V.4 will require the energy and angular momentum fluxes contributed by the second-order metric perturbation $h^{(2)}_{\alpha\beta}$. Such results are currently only available for quasicircular orbits around a Schwarzschild black hole [164, 165]. Extending that calculation to more generic binary configurations will require extending the current 2SF methods (detailed in Refs. [117, 149, 116, 156, 54]). References [155, 151, 105, 168, 148, 123, 124, 41, 71, 70, 166] describe work toward that goal.
As explained in Sec. V.4, a truly complete 1PA waveform would also require calculation of the 1PA Schott corrections $\delta P^{\rm Schott}_{A}$ to the constants of motion. Although these terms are known for quasicircular orbits [154, 80], they are not yet known for eccentric orbits. Moreover, 1PA fluxes through the primary’s horizon, as well as recently discovered “memory distortion” terms in the fluxes to infinity [54], must also be included for a genuinely complete 1PA waveform. These missing pieces merit further study, but we expect all of them to be numerically small relative to other 1PA terms.
In lieu of 2SF flux results for eccentric orbits, there is a closer-at-hand opportunity to develop 1PA waveform models: combining SF and PN information. As demonstrated in Refs. [88, 87] for quasicircular orbits, writing the waveform in terms of balance laws allows for straightforward hybridization of SF and PN information [56, 36, 51, 112, 85], and the resulting waveform model might be sufficiently accurate for high-precision EMRI science with LISA (though more work is needed to assess this, particular for more generic binary configurations), as well as being highly accurate across the full range of possible mass ratios. To see how this hybridization works, note that in both SF and PN theory, we can write the waveform in an invariant form:
$h_{+}-ih_{\times}$ $=\frac{1}{r}\sum_{\ell m}h_{\ell m}(m_{A},{\cal P}_{A},\varphi_{A}){}_{-2}Y_{\ell m}(\theta,\phi),$ (107)
$\frac{d\varphi_{A}}{ds}$ $=\Omega_{A}(m_{B},{\cal P}_{B}),$ (108)
$\frac{d{\cal P}_{A}}{ds}$ $=-{\cal F}_{A}(m_{B},{\cal P}_{B}),$ (109)
as follows from the general arguments of Sec. V.4, where we have defined $m_{A}=(m_{1},m_{2})$ and the corrected constants of motion ${\cal P}_{A}=P_{A}+\delta P^{\rm Schott}_{A}$. For completeness, we also note that the masses evolve according to the fluxes through each black hole’s horizon ${\cal H}_{A}$,
$\frac{dm_{A}}{ds}={\cal F}^{{\cal H}_{A}}_{A}(m_{B},{\cal P}_{B}),$ (110)
as do the spins (which cannot remain zero at 1PA order but which we elide here for simplicity). Hybridization can then be carried out immediately at the level of the invariant functions $\Omega_{A}(m_{B},{\cal P}_{B})$, ${\cal F}_{A}(m_{B},{\cal P}_{B})$, and $h_{\ell m}(m_{A},\varphi_{A},{\cal P}_{A})$. This is easily done using composite expansions: for a function $f(m_{A},\varphi_{A},{\cal P}_{A})$, given an expansion $f^{\rm SF}$ in powers of the mass ratio and a PN expansion $f^{\rm PN}$, we write
$f=f^{\rm SF}+f^{\rm PN}-f^{\text{SF-PN}}$ (111)
where $f^{\text{SF-PN}}$ is $f^{\rm PN}$ re-expanded in powers of the mass ratio to the same order as $f^{\rm SF}$, serving to remove otherwise doubly counted information. In the present case, one could for instance use the 3PN fluxes for non-spinning eccentric orbits [13, 12, 14, 152].
One could also calculate inputs for such a hybrid model using PM results for scattering orbits. This could be achieved using the “boundary to bound” link between scattering data and bound-orbit data [55, 95, 94, 143, 50, 78, 3, 77, 98]: the frequencies $\Omega_{A}(m_{B},{\cal P}_{B})$ and fluxes ${\cal F}_{A}(m_{B},{\cal P}_{B})$ for bound orbits can be obtained directly from the scattering angle, elapsed coordinate time, and total emitted fluxes for scattering orbits; the first two of these can also be obtained from the single relationship $J_{r}(m_{A},{\cal P}_{B})$ via the first law of of black hole binary mechanics [77]. This prospect of utilizing PM results is particularly promising because 2SF terms in the PM fluxes should soon be available [73, 72, 22, 23, 68, 24], complementing PN expansions of the 2SF fluxes. However, this approach will require extending boundary-to-bound relationships, which (in their current form) break down beyond 0SF order due to nonlocal-in-time tail effects [50, 67]; alternatively, the program embarked on in Ref. [31] can sidestep this breakdown through a separate treatment of nonlocal terms.
We expect such hybridizations, whether using only SF and PN information or SF, PN, and PM information all together, to be especially valuable in the case of a spinning, Kerr primary. 2SF flux data for eccentric orbits will be particularly challenging to generate in that case, while, as pointed out at the beginning of this section, all the other 1PA information can be readily computed within self-force theory by employing the same tools we have used in this paper [157, 121, 29, 119].
Another possible use for the 1SF constants of motion is in the construction of EOB models. The Hamiltonian in action-angle coordinates is a gauge-invariant representation of the conservative sector of the dynamics, simply equal to the energy as a function of the action variables: $H=E(J_{r},L,m_{1},m_{2})$. Such a representation was employed in the original motivating discussion for EOB [42], and one could build EOB models directly in action-angle coordinates. The invariant 1SF contribution to the Hamiltonian could then be immediately incorporated over the full range of orbital parameters. Even if such an approach is not pursued, 1SF data for generic eccentricities should offer a path to improving SF-informed EOB models for eccentric binaries, such as the TEOB model in Ref. [120], which have historically been limited to low-order small-eccentricity expansions of 1SF results [103, 7, 28].
Outside of these direct uses in waveform models, conserved quantities provide useful, invariant ways of analyzing the dynamics. For example, as originally outlined by Le Tiec in Ref. [103], 1SF definitions of $(E,L,J_{r})$ provide invariant characterizations of surfaces in phase space such as circular orbits (defined by $J_{r}=0$) and the separatrix (defined by $\Omega_{r}=0$). Using these definitions, in this paper we have computed the location of circular orbits in the $(p,e)$ plane, as an invariant curve $e_{\odot}(p)$, in the process highlighting subtleties in the definition of eccentricity for accelerated orbits. Likewise, we have computed the location of the separatrix in the $(E,L)$ plane, as an invariant curve $E_{\star}(L)$ [noting its location $p_{\star}(e)$ is unchanged from the geodesic case as a consequence of our definition of $p$ and $e$]. Another curve one might analyze, as already noted by Le Tiec, is the isofrequency degeneracy curve defined by ${\rm det}|\partial\Omega_{A}/\partial P_{B}|=0$. We leave this to future work.
Acknowledgements.
DT thanks Benjamin Leather for providing data associated to Ref. [104] and Leor Barack for insightful conversation about the ISCO shift. This work used the Black Hole Perturbation Toolkit [34], PNpedia [153] and the pybhpt package [122]. We acknowledge the support of the ERC Consolidator/UKRI Frontier Research Grant GWModels (selected by the ERC and funded by UKRI [grant number EP/Y008251/1]). AP additionally acknowledges the support of a Royal Society University Research Fellowship.
Appendix A Elliptic integrals
We recall the definitions and properties of elliptic integrals, which are used to explicitly express the geodesic frequencies in terms of $(p,e)$ in Sec. III.
We adopt the conventions of Wolfram Language when defining elliptic integral functions [173] (these differ from the definition of the NIST Digital Library of Mathematical Functions by the power of the last argument [46]). Consequently, Legendre’s complete elliptic integrals of the first, second, and third kind respectively take the forms
$\mathrm{K}(m)$ $=\int_{0}^{\pi/2}\frac{\mathrm{d}\theta}{\sqrt{1-m\sin^{2}\theta}}\,$, (112a)
$\mathrm{E}(m)$ $=\int_{0}^{\pi/2}\mathrm{d}\theta\,{\sqrt{1-m\sin^{2}\theta}}\,$, (112b)
$\Pi(n|m)$ $=\int_{0}^{\pi/2}\frac{\mathrm{d}\theta}{(1-n\sin^{2}\theta)\sqrt{1-m\sin^{2}\theta}}\,.$ (112c)
In order to perform the PN expansions of the frequencies, we will require the following small-$m$ expansions [173, 170]:
$\mathrm{K}(m)$ $=\frac{\pi}{2}\sum_{i=0}^{\infty}\frac{[(1/2)_{i}]^{2}m^{i}}{(i!)^{2}}\,$, (113a)
$\mathrm{E}(m)$ $=\frac{\pi}{2}\sum_{i=0}^{\infty}\frac{(-1/2)_{i}\,(1/2)_{i}m^{i}}{(i!)^{2}}\,$, (113b)
$\Pi(n|m)$ $=\frac{\pi}{2}\sum_{i=0}^{\infty}\frac{[(1/2)_{i}]^{2}}{(i!)^{2}}\Bigg[\frac{i!}{n^{i}\sqrt{1-n}(1/2)_{i}}$
$\qquad\qquad\qquad-\frac{2i}{n}\sum_{j=0}^{i-1}\frac{(1-i)_{j}}{(3/2)_{j}}\left(1-\frac{1}{n}\right)^{j}\Bigg]m^{i}\,$, (113c)
where $(a)_{i}=\Gamma(a+n)/\Gamma(a)=a(a+1)\cdots(a+i-1)$ is the Pochhammer symbol. In practice, Mathematica is able to directly perform the expansions of the associated functions EllipticK and EllipticE using the Series function, but this does not work with EllipticPi; it was necessary to replace the latter by its representation in terms of the Appell hypergeometric function of two variables $F_{1}(\cdots)$, denoted as AppellF1 in Mathematica. This relation reads $\Pi(n|m)=\frac{\pi}{2}F_{1}\left(\frac{1}{2};1,\frac{1}{2};1;n,m\right)$, and one can then proceed with the small-$m$ expansion of the latter using the Series function.
We also report the following relation between elliptic integrals, which is given in Eq. (17.7.17) of Ref. [2]:
$\Pi(n|m)=\frac{n(m-1)}{(n-1)(n-m)}\Pi\!\left(\!\frac{n-m}{n-1}\Big|m\!\right)-\frac{m}{n-m}K(m)\,.$ (114)
This formula allows us to reconcile different formulations for the fundamental frequencies that are present in the literature by noticing that, for m=4e/(p6+2e)m=4e/(p-6+2e), we have
$n=-\frac{2e}{1-e}$ $\Rightarrow\frac{n-m}{m-1}=\frac{2e(p-4)}{(1+e)(p-6+2e)}\,$, (115a)
$n=\frac{4e}{p+2e-2}$ $\Rightarrow\frac{n-m}{m-1}=\frac{16e}{(p-2-2e)(p-6+2e)}\,.$ (115b)
Appendix B Circular orbits from periastron advance
In this Appendix, we derive the circular link between $p$ and $e$ at postgeodesic order in terms of $\rho(x)$, and deduce the relation between $\rho(1/p)$ and $J_{r(1)}(p,0)$, which is used in the study of circular orbits in Sec. V.
First, we recall that $y=(m_{1}\Omega_{\phi})^{2/3}$ was defined in Eq. (41), and we recall the definition $x=[(m_{1}+m_{2})\Omega_{\phi}]^{2/3}$. We then find the relation
$x=y(1+\varepsilon)^{2/3}=y\left[1+\frac{2}{3}\varepsilon+\mathcal{O}(\varepsilon^{2})\right]\,.$ (116)
Following [18, 16, 19], we also define
$W=K^{-2}=\left(\frac{\hat{\Omega}_{r}}{\hat{\Omega}_{\phi}}\right)^{2}\,.$ (117)
On a circular orbit, the two frequencies are not independent; they satisfy a circular link $\hat{\Omega}_{r}=\hat{\Omega}_{r}^{\odot}(\hat{\Omega}_{\phi})$. This link can be expressed in terms of $W$, which reads [16]
$W^{\odot}=\left(\!\frac{\hat{\Omega}_{r}^{\odot}(\hat{\Omega}_{\phi})}{\hat{\Omega}_{\phi}}\!\right)^{2}$ $=1-6x+\varepsilon\rho(x)+\mathcal{O}(\varepsilon^{2})$
$=1-6y+\varepsilon\bigl[\rho(y)-4y\bigr]+\mathcal{O}(\varepsilon^{2}),$ (118)
where $\rho(x)$ is given (i) numerically in Table II of Ref. [16] for $x\in[1/80,1/6]$ and in the Supplemental Material of [159] for $x\in[1/1000,1/10]$; and (ii) as a PN series in Eq. (9) of Ref. [30].
At geodesic order, $p_{\odot}=1/y+\mathcal{O}(\varepsilon)$ and $e_{\odot}^{2}=\mathcal{O}(\varepsilon)$, so we introduce the ansatz
$p_{\odot}$ $=1/y+\varepsilon\,\delta p_{\odot}+\mathcal{O}(\varepsilon^{2}),$ (119)
$e_{\odot}^{2}$ $=\varepsilon\,\delta e_{\odot}^{2}+\mathcal{O}(\varepsilon^{2}).$ (120)
We plug this ansatz into the exact expression of $(\hat{\Omega}_{\phi},\hat{\Omega}_{r})$ in terms of $(p,e)$ and Taylor expand in $\varepsilon$; see App. A for how to do this in practice. Thanks to these expressions, we find that
$y=\hat{\Omega}_{\phi}^{2/3}(p_{\odot},e_{\odot}^{2})$ $=y+\varepsilon\mathcal{A}(y,\delta p,\delta e),$ (121a)
$W^{\odot}=K^{-2}(p_{\odot},e_{\odot}^{2})$ $=1-6y+\varepsilon\,\mathcal{B}(y,\delta p,\delta e),$ (121b)
where we have ignored $\mathcal{O}(\varepsilon^{2})$ terms and where
$\mathcal{A}(y,\delta p,\delta e)$ $=-y^{2}\delta p+\frac{y(-1+10y-22y^{2})}{1-8y+12y^{2}}\delta e_{\odot}^{2}\,$, (122a)
$\mathcal{B}(y,\delta p,\delta e)$ $=6y^{2}\delta p+\frac{3y^{2}}{-2+12y}\delta e_{\odot}^{2}\,.$ (122b)
Equating this to (B), we obtain the constraints
$\mathcal{A}(y,\delta p,\delta e)$ $=0\,$,\, (123a)
$\mathcal{B}(y,\delta p,\delta e)$ $=\rho(y)-4y\,$, (123b)
which translate to
$\delta p_{\odot}$ $=-\frac{2(1-10y+22y^{2})\big[4y-\rho(y)\big]}{3y^{2}(4-39y+86y^{2})}\,$, (124a)
$\delta e_{\odot}^{2}$ $=\frac{2(1-2y)(1-6y)\big[4y-\rho(y)\big]}{3y(4-39y+86y^{2})}\,.$ (124b)
Thus, using $y=1/p+\mathcal{O}(\varepsilon)$, we have
$\delta e_{\odot}^{2}(p)=\frac{2(p-6)(p-2)}{3(86-39p+4p^{2})}\bigl[4-p\,\rho(1/p)\bigr]\,.$ (125)
Injecting this expression into (62) finally leads to
$J_{r(1)}^{\odot}(p)=\frac{p^{3/2}(p-6)^{3/2}\bigl[p\,\rho(1/p)-4\bigr]}{3\sqrt{p-3}(86-39p+4p^{2})}\,.$ (126)
Using our PN expansion for $J_{r}^{(1)}(p,0)$ and the PN expansion of $\rho(x)$ provided in Ref. [30], we have checked that the previous result is consistent through 9PN order.
Appendix C Numerical evaluation of circular limits
We describe our numerical approach to computing 1SF corrections along circular orbits and on the ISCO, as reported in Secs. IV.3 and V.3. While frequency derivatives are well defined in the circular limit, the Jacobian elements $\partial e/\partial\hat{\Omega}_{r}$ and $\partial e/\partial\hat{\Omega}_{\phi}$ are singular at $e=0$. Thus, care must be taken when numerically evaluating derivatives in this limit. First, we expand all quantities about $e=0$ and find
$\frac{\partial p}{\partial\hat{\Omega}_{r}}$ $=\frac{4p^{3}\sqrt{p-6}(p^{2}-10p+22)}{3(4p^{2}-39p+86)}+\mathcal{O}(e^{2}),$ (127a)
$\frac{\partial p}{\partial\hat{\Omega}_{\phi}}$ $=-\frac{2p^{5/2}(2p^{3}-32p^{2}+165p-266)}{3(4p^{2}-39p+86)}+\mathcal{O}(e^{2}),$ (127b)
$\frac{\partial e}{\partial\hat{\Omega}_{r}}$ $=-\frac{2p^{2}(p-6)^{3/2}(p-2)}{3(4p^{2}-39p+86)e}+\mathcal{O}(e),$ (127c)
$\frac{\partial e}{\partial\hat{\Omega}_{\phi}}$ $=\frac{2p^{3/2}(p-6)(p-2)(p-8)}{3(4p^{2}-39p+86)e}+\mathcal{O}(e).$ (127d)
We then expand the redshift in small eccentricity:
$\langle z_{(1)}\rangle(p,e)=z_{(1)}^{[0]}(p)+e^{2}\,z^{[2]}_{(1)}(p)+\mathcal{O}(e^{4})\,$, (128)
where $z_{(1)}^{[0]}(p)=\langle z_{(1)}\rangle(p,0)$ and
$z^{[2]}_{(1)}(p)=\frac{1}{2}\left.\frac{\partial^{2}\langle z_{(1)}\rangle}{\partial e^{2}}\right|_{e=0}\,.$ (129)
In Eq. (128), we have leveraged ${\partial\langle z_{(1)}\rangle}/{\partial e}=0$ for $e=0$. Note that, for this work, we directly compute $z_{(1)}^{[0]}(p)$, but obtain $z_{(1)}^{[2]}(p)$ via finite differencing.
The 1SF corrections to the constants of motion can then be expressed in terms of radial derivatives of the redshift at zero eccentricity and the eccentricity corrections $z^{[2]}_{(1)}$ (or, alternatively, first-derivatives of the redshift with respect to $e^{2}$), leading to
$\hat{E}_{(1)}^{\odot}(p)$ $=\frac{1}{2}z_{(1)}^{[0]}(p)+\frac{p(p-2)}{3(4p^{2}-39p+86)}\ \frac{dz_{(1)}^{[0]}}{dp}$
$\quad+\frac{4(p-6)(p-2)}{3(4p^{2}-39p+86)}\ z^{[2]}_{(1)}(p)\,$, (130a)
$\hat{L}_{(1)}^{\odot}(p)$ $=\frac{p^{5/2}(2p^{3}-32p^{2}+165p-266)}{3(4p^{2}-39p+86)}\ \frac{dz_{(1)}^{[0]}}{dp}$
$\quad-\frac{2p^{3/2}(p-6)(p-2)(p-8)}{3(4p^{2}-39p+86)}\ z^{[2]}_{(1)}(p)\,$, (130b)
$\hat{J}_{r(1)}^{\odot}(p)$ $=-\frac{2p^{3}\sqrt{p-6}(p^{2}-10p+22)}{3(4p^{2}-39p+86)}\ \frac{dz_{(1)}^{[0]}}{dp}$
$\quad+\frac{2p^{2}{(p-6)^{3/2}}(p-2)}{3(4p^{2}-39p+86)}\ z^{[2]}_{(1)}(p)\,.$ (130c)
Further simplifications need to be made when approaching the ISCO. Expanding in $\delta_{0}\equiv\delta|_{e=0}=p-6$ leads to the near-ISCO ansatz
$z^{[0]}_{(1)}(p)$ $=a_{0}+a_{1}\delta_{0}+a_{2}\delta_{0}^{2}+\mathcal{O}(\delta_{0}^{3}),$ (131a)
$z^{[2]}_{(1)}(p)$ $=b_{-1}\delta_{0}^{-1}+b_{0}+b_{1}\delta_{0}+\mathcal{O}(\delta_{0}^{2}),$ (131b)
$z^{[0]}_{(1)}(p)$ $=a_{0}+a_{1}(p-6)+a_{2}(p-6)^{2}+\mathcal{O}\left[(p-6)^{3}\right],$ (132a)
$z^{[2]}_{(1)}(p)$ $=b_{-1}\delta_{0}^{-1}+b_{0}+b_{1}(p-6)+\mathcal{O}\left[(p-6)^{2}\right],$ (132b)
where the coefficients $a_{n}$ and $b_{n}$ are not known exactly. In this work, we compute these coefficients by fitting the expansions in (132) to numerical values of $z^{[0]}_{(1)}(p)$ and $z^{[2]}_{(1)}(p)$ near the ISCO. Inserting (132) into our expression for the radial action (130), we find
$\hat{J}_{r(1)}^{\mathrm{ISCO}}$ $=-24(b_{-1}+3a_{1})\sqrt{\delta_{0}}$ (133)
$\quad-2(34b_{1}+12b_{0}+63a_{1}+72a_{2})\delta_{0}^{3/2}+\mathcal{O}(\delta_{0}^{5/2})\,.$
Recalling from Eq. (63) that $\hat{J}_{r(1)}^{\odot}(p)/(p-6)^{3/2}$ must be regular and finite as $p\rightarrow 6$, we conclude that $b_{-1}=-3a_{1}$; thus, the first term vanishes. This conclusion is consistent with our numerics. Consequently, at leading order we have
$\hat{E}_{(1)}^{\mathrm{ISCO}}$ $=\frac{1}{2}a_{0}+2a_{1},$ (134a)
$\hat{L}_{(1)}^{\mathrm{ISCO}}$ $=12\sqrt{6}a_{1},$ (134b)
$\hat{J}_{r(1)}^{\mathrm{ISCO}}$ $=2(39a_{1}-72a_{2}-12b_{0})(p-6)^{3/2},$ (134c)
with the post-geodesic correction to the radial action vanishing exactly on the ISCO. Note that these expressions are consistent with the relations in Ref. [99] between the redshift, the binding energy, and the angular momentum in the circular limit, once we have re-expanded in $x=[(m_{1}+m_{2})\Omega_{\phi}]^{2/3}$ in terms of $p$ to 1SF order.
From this, we can also express $\rho(1/6)$ and $\Omega^{\mathrm{ISCO}}_{(1)}$ in terms of these expansion terms,
$\rho(1/6)$ $=\frac{2}{3}+\sqrt{2}(4b_{0}-13a_{1}+24a_{2}),$ (135a)
$\frac{\Omega^{\mathrm{ISCO}}_{(1)}}{\Omega^{\mathrm{ISCO}}_{(0)}}$ $=\frac{3}{\sqrt{2}}(4b_{0}-13a_{1}+24a_{2}).$ (135b)
By fitting near-ISCO redshift values, we obtain the following numerical estimates for the expansion coefficients:
$a_{0}\;$ $=\phantom{-}0.14801375(1)\,$,
$a_{1}\;$ $=-0.027291088(3)\,$,
$a_{2}\;$ $=\phantom{-}0.0073291(1)\,$, (136)
$b_{-1}\;$ $=\phantom{-}0.081873(2)\,$,
$b_{0}\;$ $=-0.10309(3)\,.$
We obtain these results using a least squares fit with scipy.optimize.curve_fit within the Python SciPy library [162]. We fit to expansions with the functional forms
$f(\delta_{0})=\sum_{n=n_{0}}^{N}A_{n}\delta_{0}^{n},$ (137)
where $n_{0}=-1$ for $z^{[2]}_{(1)}(p)$ and $\rho(1/p)$ and $n_{0}=0$ for all other quantities. The fit is weighted by the estimated numerical errors of the input data. We produce a family of fits by varying the number of coefficients $A_{n}$ and the range of input grid data that we fit over. In particular, we vary $4\leq N-n_{0}\leq 16$ and grid values in the range $\delta_{\mathrm{min}}\leq\delta_{0}\leq 5$, though we do not include data from grid points with estimated relative errors >100>100. We take $5\times\sigma_{nn}$ to be the estimated error in the fit parameter $A_{n}$, where $\sigma_{nm}^{2}$ is the covariance matrix produced from the fit. From the set of fits, we select the three fits that produce the smallest estimated errors $5\times\sigma_{n_{0}n_{0}}$ for $A_{n_{0}}$. If the variance of these three estimated values of $A_{n_{0}}$ is smaller than their estimated errors, then we use the fit with the smallest error. If the variance is greater, than we take the mean and variance of these three fits to get the values and errors of the fit coefficients $A_{n}$.
Appendix D Corrections to the separatrix curve
[…]
which leads to the relations
[…]
Inserting these into Eq. (79b), we find
$\hat{E}^{\star}_{(1)}(\hat{L})=E^{\star}_{(1)}(e_{(0)}^{\star}(\hat{L}))-\left[\frac{\mathrm{d}\hat{E}^{\star}_{(0)}/\mathrm{d}e}{\mathrm{d}\hat{L}^{\star}_{(0)}/\mathrm{d}e}\hat{L}^{\star}_{(1)}\right]_{e=e_{(0)}^{\star}(\hat{L})}$, (140)
where $\mathrm{d}\hat{E}^{\star}_{(0)}/\mathrm{d}e$ and $\mathrm{d}\hat{L}^{\star}_{(0)}/\mathrm{d}e$ can be computed from Eqs. (78a) and (78b), respectively.
[…]
$\pi_{A}(\hat{P}_{B},\epsilon)=\pi_{A}^{(0)}(\hat{P}_{B})+\varepsilon\pi_{A}^{(1)}(\hat{P}_{B}),$ (141)
with
$\pi_{A}^{(1)}(\hat{P}_{B})=-\frac{\partial\pi_{A}^{(0)}}{\partial\hat{P}_{B}}\hat{P}_{B}^{(1)}.$ (142)
Demanding the $p=6+2e$ on the separatrix leads to the relation
$p^{\star}_{(0)}(\hat{P}_{B})+\varepsilon p^{\star}_{(1)}(\hat{P}_{B})=6+2e^{\star}_{(0)}(\hat{P}_{B})+\varepsilon\,2e^{\star}_{(1)}(\hat{P}_{B}).$ (143)
[…]
$\hat{E}^{\star}_{(1)}(\hat{L})=\frac{p^{\star}_{(1)}(\hat{E}^{\star}_{(0)}(\hat{L}),\hat{L})-2e^{\star}_{(1)}(\hat{E}^{\star}_{(0)}(\hat{L}),\hat{L})}{2\partial_{E}e^{\star}_{(0)}(\hat{E}^{\star}_{(0)}(\hat{L}),\hat{L})-\partial_{E}p^{\star}_{(0)}(\hat{E}^{\star}_{(0)}(\hat{L}),\hat{L})}$. (144)
[…]
with the first-order correction given by Eq. (46). Inserting Eq. (84) into the expression above, and demanding that $\hat{\Omega}_{r}$ vanishes on the separatrix, we then find
$\hat{E}^{\star}_{(1)}(\hat{L})=-\frac{\hat{\Omega}_{r(1)}(\hat{E}^{\star}_{(0)}(\hat{L}),\hat{L})}{\partial_{\hat{E}}\hat{\Omega}_{r(0)}(\hat{E}^{\star}_{(0)}(\hat{L}),\hat{L})}$. (146)
Appendix E Relations between partial derivatives
Here, we derive relations between (i) the partial derivatives of the conserved energy, angular momentum, and radial action with respect to $p$ and $e$, and (ii) the second-order partial derivatives of the redshift $\langle z_{1}\rangle$ with respect to $p$ and $e$. These formulas are needed in one of the formulations f for waveform generation described in Sec. V.4.
We know the explicit expressions for $\hat{\Omega}_{A}(\pi_{B})$, but not for the inverses $\pi_{A}(\hat{\Omega}_{B})$. However, it is straightforward to compute, by matrix inversion, the first partial derivatives of $(p,e)$ with respect to $(\Omega_{r},\Omega_{\phi})$ as functions of $(p,e)$; see Eq. (105). Here, we extend this method to the case of second derivatives with respect to $p$ and $e$.
Using the chain rule, we first obtain the identity
$0=\frac{\partial^{2}\pi_{A}}{\partial\pi_{B}\partial\pi_{C}}=\frac{\partial\hat{\Omega}_{D}}{\partial\pi_{C}}\frac{\partial\hat{\Omega}_{E}}{\partial\pi_{B}}\frac{\partial^{2}\pi_{A}}{\partial\hat{\Omega}_{D}\partial\hat{\Omega}_{E}}+\frac{\partial\pi_{A}}{\partial\hat{\Omega}_{E}}\frac{\partial^{2}\hat{\Omega}_{E}}{\partial\pi_{C}\partial\pi_{B}}.$ (147)
Contracting both sides with $(\partial\pi_{C}/\partial\hat{\Omega}_{D})\times(\partial\pi_{B}/\partial\hat{\Omega}_{E})$, we find that
$\frac{\partial^{2}\pi_{A}}{\partial\hat{\Omega}_{B}\partial\hat{\Omega}_{C}}=-\frac{\partial\pi_{A}}{\partial\hat{\Omega}_{F}}\frac{\partial\pi_{D}}{\partial\hat{\Omega}_{B}}\frac{\partial\pi_{E}}{\partial\hat{\Omega}_{C}}\frac{\partial^{2}\hat{\Omega}_{F}}{\partial\pi_{D}\partial\pi_{E}}.$ (148)
For any function $f(p,e)$, we can use this identity to obtain
$\frac{\partial^{2}f}{\partial\hat{\Omega}_{B}\partial\Omega_{C}}=\frac{\partial\pi_{D}}{\partial\hat{\Omega}_{B}}\frac{\partial\pi_{E}}{\partial\hat{\Omega}_{C}}\Bigg(\frac{\partial^{2}f}{\partial\pi_{D}\partial\pi_{E}}-\frac{\partial f}{\partial\pi_{A}}\frac{\partial\pi_{A}}{\partial\hat{\Omega}_{F}}\frac{\partial^{2}\hat{\Omega}_{F}}{\partial\pi_{D}\partial\pi_{E}}\Bigg).$ (149)
The 1SF correction to the energy, angular momentum and radial action are given in terms of derivatives $\partial\langle z_{(1)}\rangle/\partial\hat{\Omega}_{A}$ in Eq. (27). Using the identity (149) with $f=\langle z_{(1)}\rangle$, we can write their derivatives with respect to $(p,e)$ as
$\frac{\partial\hat{E}_{(1)}}{\partial\pi_{A}}$ $=-\frac{\hat{\Omega}_{C}}{2}\frac{\partial\pi_{B}}{\partial\hat{\Omega}_{C}}\Bigg(\frac{\partial^{2}\langle z_{(1)}\rangle}{\partial\pi_{A}\partial\pi_{B}}-\frac{\partial\langle z_{(1)}\rangle}{\partial\pi_{D}}\frac{\partial\pi_{D}}{\partial\hat{\Omega}_{E}}\frac{\partial^{2}\hat{\Omega}_{E}}{\partial\pi_{A}\partial\pi_{B}}\Bigg)\,$, (150a)
$\frac{\partial\hat{L}_{(1)}}{\partial\pi_{A}}$ $=-\frac{1}{2}\frac{\partial\pi_{B}}{\partial\hat{\Omega}_{\phi}}\Bigg(\frac{\partial^{2}\langle z_{(1)}\rangle}{\partial\pi_{A}\partial\pi_{B}}-\frac{\partial\langle z_{(1)}\rangle}{\partial\pi_{D}}\frac{\partial\pi_{D}}{\partial\hat{\Omega}_{E}}\frac{\partial^{2}\hat{\Omega}_{E}}{\partial\pi_{A}\partial\pi_{B}}\Bigg)\,$, (150b)
$\frac{\partial\hat{J}_{r(1)}}{\partial\pi_{A}}$ $=-\frac{1}{2}\frac{\partial\pi_{B}}{\partial\hat{\Omega}_{r}}\Bigg(\frac{\partial^{2}\langle z_{(1)}\rangle}{\partial\pi_{A}\partial\pi_{B}}-\frac{\partial\langle z_{(1)}\rangle}{\partial\pi_{D}}\frac{\partial\pi_{D}}{\partial\hat{\Omega}_{E}}\frac{\partial^{2}\hat{\Omega}_{E}}{\partial\pi_{A}\partial\pi_{B}}\Bigg)\,.$ (150c)